### v2

# Image Inpainting based on Visual-Neural-Inspired Specific Object-of-Interest Imaging Technology

Abstract
Conventional image inpainting methods operate on whole images. Drawing on cortical processing principles this study addresses the bottlenecks of conventional holistic image inpainting methods–susceptibility to informational redundancy and low computational efficiency under occlusions and complex backgrounds, by proposing a novel framework: ”Specific Object-of-Interest Imaging”. This stage extracts and encodes object-level representations from complex scenes, generating semantic and structural priors that can be seamlessly integrated into any inpainting framework. On our Teapot object dataset, Elephant object dataset, Giraffe object dataset and Zebra object dataset experimental validation demonstrates–compared with the repair model without using this method, the repair model using this method has more advantages, across metrics including SSIM, PSNR, MAE, and LPIPS, while maintaining robustness in extreme scenarios (low illumination, high noise, multi-object occlusion, motion blur). Theoretical analysis integrated with cognitive neuroscience perspectives reveals profound correlations between the ”object precedence perception” mechanism and dynamic feature modulation in visual cortices (V1–V4). We demonstrate that incorporating Stage I significantly enhances object consistency and semantic coherence across various mainstream inpainting models. This work not only achieves efficient and precise target-centric imaging but also pioneers interdisciplinary pathways bridging brain-inspired computational frameworks with advanced image inpainting techniques.
keywords:
Image Inpainting , Specific Object-of-Interest Imaging , Visual Neural Mechanisms , Deep Neural Network
journal: Neural Networks
orrefcempty
[…]
organization=School of Electronic Science and Engineering (School of Microelectronics), South China Normal University, addressline=Foshan, 528000, country=China
[…]
organization=St. Petersburg School of Shipbuilding and Ocean Technology, Guangdong Ocean University, addressline=Zhanjiang 524088, country=China
[…]
organization=Robotics Laboratory, Institute of Automation and Control Processes, addressline=Far Eastern Branch of the Russian Academy of Sciences, city=Vladivostok, postcode=690041, country=Russia
[…]
organization=Robotic Laboratory, Institute of Automation and Control Processes FEB RAS, city=Vladivostok, postcode=690041, country=Russia
{graphicalabstract}
{highlights}
Research highlight 1
Research highlight 2
1 Introduction
Image Inpainting, a foundational and critically important task in computer vision and image processing, aims to restore missing or corrupted regions. Such impairments may arise from accidental occlusions during capture, sensor-induced noise, data loss in encoding/decoding or transmission, intentional object removal in post-processing, or physical deterioration in historical photographs. The capacity to restore image integrity, visual authenticity, and semantic plausibility renders this technology indispensable across diverse applications, including cultural heritage preservation, digital photography enhancement, medical image reconstruction, cinematic visual effects production, and the interpretation of occluded scenarios in autonomous driving systems. However, producing highly accurate and semantically coherent inpainting results remains a fundamental challenge.
Conventional inpainting methods primarily rely on intrinsic statistical properties and local structural information of images [1, 2, 3, 4, 5, 6, 7, 8]. Canonical approaches include diffusion- and patch-based algorithms, whose core principle involves leveraging neighborhood information from intact regions to fill missing areas via propagation or copy-paste mechanisms. While adequate for small-scale, texturally regular regions, these methods exhibit significant limitations when confronted with extensive missing areas, complex structural boundaries, or semantically rich regions. They often generate artifacts such as blurred textures, fractured structural continuity, and semantically inconsistent content that undermines global image context. These limitations significantly constrain their effectiveness in real-world, high-fidelity imaging applications.
The rapid development of deep learning—particularly Convolutional Neural Networks (CNNs), Generative Adversarial Networks (GANs), and emerging Transformer- and Diffusion-based models—has driven breakthroughs in complex pattern learning and high-fidelity image synthesis, thereby catalyzing a paradigm shift in image inpainting [9, 10, 11, 12, 13, 14, 15]. These data-driven approaches capture rich image priors and deep semantic representations from large-scale datasets, thereby enabling models to predict missing content with improved accuracy. These models reconstruct intricate textural details while maintaining plausibility. For example, restored objects exhibit geometrically consistent structures, physically coherent illumination, and contextually appropriate depth relationships. State-of-the-art models can generate diverse yet equally plausible results, demonstrating performance that substantially surpasses conventional methods.
Despite these advances and their widespread deployment, researchers still face persistent and critical challenges. A pivotal yet often overlooked issue is the inefficiency and inaccuracy stemming from informational redundancy. In many real-world applications, inpainting is fundamentally object-centric rather than indiscriminately processing the entire frame. Under such conditions, prevailing whole-image inpainting models — encompassing nearly all contemporary deep learning methods — incur substantial computational overhead by processing redundant background information. For instance, autonomous vehicles equipped with 8K cameras can generate up to 1.5 TB/s of data, with nearly 90% consisting of redundant background or irrelevant content; military reconnaissance satellites transmit massive raw imagery, with less than 0.1% containing actionable intelligence; and in medical X-ray imaging and industrial non-destructive testing, conventional area-array detectors achieve less than 0.13% efficiency in extracting useful information. Whole-image models consume substantial computational resources analyzing redundant background information, leading to severe resource wastage and reduced accuracy in target-object reconstruction due to background interference, calling for a paradigm shift from holistic processing to targeted, object-focused inference.
In confronting these challenges, biological vision systems offer an elegant and efficient solution. A core component of biological vision is the attention mechanism, which prioritizes task-relevant targets while suppressing redundant background information. This enables efficient, accurate, and energy-optimized information processing for target recognition and imaging in cluttered or dynamic scenes. This “object-first perception” strategy offers a promising solution for addressing low information density and background noise in inpainting: an inpainting pipeline could first localize and model target-object structures while excluding irrelevant background as interference. Such an approach would enhance imaging focus, computational efficiency, and output quality.
Building on this analysis of the current state of inpainting, its technical bottlenecks, and biological inspiration, we propose a novel front-end paradigm: Image Inpainting via Visual-Neural-Inspired Specific Object-of-Interest Imaging technology(SIOI). Inspired by biological “object-first perception”, we propose a paradigm shift: Specific Object-of-Interest Imaging (SIOI). Unlike traditional methods, SIOI decouples the ‘what’ and ‘where’ of the object from the ‘how’ of its inpainting. It acts as a generic front-end that extracts a clean object representation, suppressing background noise, and can be plugged into any existing inpainting model. Our framework decomposes the imaging process into two distinct stages. Stage 1 focuses exclusively on extracting and modeling the structural information of target objects. This stage constitutes a dedicated information-extraction process designed to maximize target-relevant features while minimizing or eliminating background interference. Stage 2 performs coarse structural restoration guided by refined priors to support detail synthesis, refines the restoration, ensuring visual coherence and semantic consistency in the final output. The key innovation lies in Stage I; Stage II is deliberately kept simple in our experiments to prove the concept. By frontloading SIOI, the pipeline significantly reduces background distractions and concentrates structural modeling on critical target regions. This substantially reduces the data volume for processing, thereby improving computational efficiency while enhancing structural continuity and semantic accuracy. As shown in Tab. 1, unlike other image-restoration methods, this bio-inspired approach aligns with the hierarchical modeling of the visual cortex in neural networks, ensuring that computational efficiency parallels human attentional selection.
Our principal contributions are:
  1. 1.
    We introduce Specific Object-of-Interest Imaging (SIOI), a novel front-end paradigm for object-centric image processing. SIOI explicitly decouples target-object extraction from texture synthesis, providing a generic mechanism to reduce background interference and generate powerful structural priors for any subsequent task.
  2. 2.
    We conduct a preliminary exploration of correlations between human vision and deep networks, integrating neuroscientific and psychological insights on hierarchical visual cortex processing to establish theoretical foundations.
  3. 3.
    We demonstrate that SIOI is an effective, plug-and-play front-end module. Through extensive experiments, we show that integrating SIOI consistently enhances the performance of various mainstream inpainting models across multiple metrics and datasets, while significantly improving robustness in extreme scenarios.
Table 1: Theoretical Framework Comparison
Model Theoretical Framework Core Mechanism Processing Unit Biological Plausibility
Traditional Digital image processing and classical machine learning techniques Handcrafted features and algorithms Entire image Not considered
Deep-Learning Data-driven approaches enabled by deep neural networks Learned feature representations Entire image Not considered
Ours Bio-inspired two-stage framework Object perception → Object inpainting Specific Object-of-Interest V1-V4 hierarchical simulation
2 Related Works
Image inpainting is a long-studied and widely applied problem whose methods have undergone substantial paradigm shifts. From early mathematical models to modern deep generative architectures. While these approaches differ in implementation, they overwhelmingly share a whole-image processing paradigm.
2.1 Traditional Mathematical Algorithms: Digital Image Processing and Machine Learning
This period emphasized leveraging low-level statistical properties, local geometric structures, and hand-crafted mathematical models. Representative approaches include: 1) PDE- and diffusion-based models [1, 16, 17], which use isophote directions or engineered diffusion tensors to propagate color and gradient information from intact regions into missing areas while preserving edge coherence; and 2) patch-based methods [18, 19, 20], which exploit non-local self-similarity by searching for patches in intact regions that best match neighborhoods around missing regions and iteratively copying them to fill holes. Traditional mathematical approaches [21, 22, 23] reconstruct missing areas by propagating local structures or copying similar patches. They perform reasonably on small, regular textures but break down under large occlusions or semantically rich regions, since they lack object-level understanding.
2.2 Data-Driven Neural Network Approaches
Deep-learning approaches substantially improve structural coherence and realism by learning high-level priors. Predominant paradigms include: (1) CNN encoder–decoder architectures [24, 4, 25, 26, 27], which exploit multi-level encoder features to progressively generate missing content in decoders, attention mechanisms and gated convolutions further improve focus on relevant regions; (2) GAN-based methods [28, 29, 30, 31, 32, 33, 34], which employ adversarial training between generators and discriminators to produce highly realistic outputs and excel at fine texture synthesis and alignment with natural image distributions; (3) Transformer-based approaches [4, 35, 36, 37, 38, 39, 40], which leverage self- and cross-attention to model long-range dependencies between patches and thereby capture global semantics and structural consistency, improving performance on complex scenes and large corruptions; and (4) diffusion models [41, 42, 43], which cast generation as a progressive denoising process that iteratively recovers image details from noise, offering fine-grained detail control, diverse outputs, and high fidelity.
Yet, most still operate on the entire image, wasting computation on irrelevant background and suffering from interference when target and background overlap. Crucially, they rarely encode explicit object guidance, limiting their ability to ensure object consistency in inpainting. Whole-image models thus suffer from several issues in object-centric tasks: 1) information redundancy and computational inefficiency — processing the entire image wastes computation on irrelevant background content, exacerbating the low effective information density typical in domains such as autonomous driving and satellite imaging; 2) background interference — complex or misleading backgrounds can corrupt the reconstruction of target object structure and semantics, especially when target and background share similar textures or colors; and 3) lack of target-specific guidance and focus — current methods rarely model explicitly which object should be targeted and therefore struggle to separate target signals from background noise during inpainting. Although some methods adopt user sketches or bounding boxes as guidance, these require extra interaction and still struggle to achieve precise target separation in complex scenes.
Thus, a biologically inspired, target-focused imaging paradigm — one that actively suppresses irrelevant background information — is imperative for object-centric tasks. Our proposed method, Inpainting via Specific Object-of-Interest Imaging , directly addresses this gap. It introduces a front-loaded Specific Object-of-Interest Imaging module that algorithmically maximizes target-relevant information while suppressing background interference. This yields highly refined, target-centric signals for subsequent inpainting stages and thereby mitigates information redundancy and background interference.
2.3 Insights from Human Visual Mechanisms for Image imaging
Psychologists and neuroscientists have reported direct correlations between deep neural networks (DNNs) and hierarchical processing in the visual cortex [44, 45]. Human vision follows a hierarchical pathway: retina $\rightarrow$ lateral geniculate nucleus (LGN) $\rightarrow$ V1 (edge extraction) $\rightarrow$ V4 $\rightarrow$ IT (semantic processing) [46, 47]. This observation motivates multi-stage imaging networks [48, 49], in which early convolutional layers focus on local structure repair (analogous to V1 edge detection), while deeper layers integrate object identity and scene context (analogous to IT semantic processing); skip connections can be viewed as a crude analogue of cortical feedback. Predictive-coding theory [50, 51] proposes that the brain leverages prior knowledge (e.g., Gestalt principles) to generate predictions and correct sensory errors. This idea parallels the generator–discriminator framework in deep inpainting [24, 4]: generators predict missing content by learning data distributions, while discriminators enforce semantic plausibility via adversarial training [52]. At the level of attentional regulation, bottom-up attention captures salient features (e.g., high-frequency details) via thalamic structures such as the pulvinar [53, 54]; this motivates algorithmic mechanisms such as saliency-map-weighted losses that emphasize abrupt edge regions [55]. Conversely, top-down attention involves prefrontal-cortex (PFC) modulation of visual responses according to task goals [54]. This inspires task-oriented strategies in models: for example, conditional vectors can steer generation to enhance discriminative feature repair (e.g., facial keypoint reconstruction) and to improve global structural consistency for scene reconstruction. Examples of biologically inspired techniques include contour interpolation via edge-consistency losses [56], motivated by V1/V2 responses to illusory contours [57, 58, 59]. Semantic guidance has been realized by mapping IT-like object activations. For instance, CLIP-guided diffusion has been used to align textual concepts with image generation [42]. Memory-augmented imaging methods simulate hippocampal–cortical episodic recall by retrieving reference segments from external memory banks [60]. Consequently, advancing image inpainting benefits from drawing principled inspiration from human visual mechanisms to identify promising directions for improvement. This ‘object-first’ perception allows efficient suppression of background noise and allocation of resources to relevant targets. Such mechanisms directly motivate our proposed Specific Object-of-Interest Imaging, which introduces an analogous front-end object-first module into the inpainting pipeline.”
In summary, while existing methods have advanced in realism and fidelity, they all share a whole-image bias that is ill-suited for object-centric tasks. Inspired by biological vision, we propose a new front-end paradigm—Specific Object-of-Interest Imaging—to fill this critical gap.
3 Proposed Methodology
This section delineates our proposed framework for Image Inpainting. Our approach introduces a novel paradigm for this task. On a conceptual level, it shares a similar goal—focusing on a specific object region—with other imaging techniques, such as [61, 62, 63, 64, 65], but the underlying methodology is uniquely our own work. The processing pipeline is illustrated in Fig. 1.
Figure 1: Specific Object-of-Interest Imaging technology.
3.1 Specific Object-of-Interest Imaging
Specific Object-of-Interest Imaging is also known as Attention-Targeted Imaging. Desimone and Duncan’s competitive-bias model [66] posits that multiple objects compete for activation resources in the visual cortex. Attention serves as a bias signal that preferentially amplifies target objects and suppresses background information. This provides a neuroscientific foundation for implementing background suppression and object-enhancement saliency selection in computational networks. O’Craven et al. [67] demonstrated with fMRI that objects act as fundamental attentional units, exhibiting holistic enhancement even when component features are spatially segregated. Other studies [68, 69] further showed that object boundaries automatically attract attention, recruiting V1–V4 regions and intensifying internal feature encoding. Perry et al. [70] reported that foveal cortical magnification increases encoding density, and that task-dependent plasticity enables high-resolution perception of behaviorally relevant regions. Carrasco and Barbot [71, 72] systematically characterized attention’s dual regulation of spatial resolution: exogenous attention enhances frequency sensitivity at fixation loci, whereas endogenous attention reallocates resolution according to task demands. Barbot and Carrasco further showed that attentional guidance amplifies high-frequency recognition, providing a basis for region-adaptive reconstruction. Ekman et al. [73] demonstrated that attentional signals propagate along object boundaries in V1, inducing internal feature enhancement — a phenomenon translatable to frequency-domain focusing mechanisms in networks (e.g., concentrating computational resources using object masks).Collectively, these findings reveal an object-based prioritization mechanism: the visual system preferentially allocates resources to enhance detailed representation of target objects. These observations directly inspired our object-detection–guided saliency-enhancement branch, which applies high-frequency component weighting to target regions. We propose an approach that integrates saliency modeling with biologically inspired, object-guided image imaging to construct a cognitively plausible, interpretable, and task-relevant image inpainting system (see Fig. 2).
Figure 2: Image inpainting framework based on the visual neural Specific Object-of-Interest Imaging.
Neuroanatomically, visual processing follows a hierarchical architecture: local feature detection (retina) $\rightarrow$ spatial integration (LGN) $\rightarrow$ orientation selectivity (V1) $\rightarrow$ holistic object representation (IT cortex). This multi-stage processing implies that a natural-scene image $k$ can be viewed as a superposition of multiple latent target objects, with attentional modulation in V4 amplifying target-specific features while suppressing background information. Grounded in neuronal population-coding theory and Hebbian plasticity, we formalize this composition as a linear superposition model:
$k=\sum_{i=1}^{M}{\{}k_{i}\}$ (1)
where $k\in\mathbb{R}^{H\times W\times C}$ denote the scene tensor, $k_{i}\in\mathbb{R}^{h_{i}\times w_{i}\times C}$ the $i$-th target object, and $M\in\mathbb{Z}^{+}$ the total number of objects. Each $k_{i}$ corresponds to a semantically complete entity in the scene (e.g., pedestrians or vehicles). Amodal content completion emphasizes restoring full object structures in occluded regions. The core objective is anatomically precise object separation: accurately extracting specific targets $k_{i}$ ($i\in\{1,\dots,M\}$) from complex backgrounds. This parallels the “cocktail-party” effect in human perception when attending to a target in cluttered scenes.
To overcome the representational fragility of traditional methods in noisy environments, we propose a biologically inspired dynamic feature-modulation paradigm. Its mathematical core is a target-sensitive measurement matrix $\Gamma$. Via a differentiable learning mechanism, $\Gamma$ induces a cross-scale feature set $\{\xi_{j}\}_{j=1}^{N}$, reflecting the hierarchical abstraction from V1 simple cells (edge/texture responses) to IT complex cells (holistic shape responses). Below we make these notions precise.
For a target index $\mu\in\{1,\dots,M\}$, we write
$\Gamma k_{\mu}\;=\;\sum_{j=1}^{N}\omega_{j,\mu}\,\xi_{j},$ (2)
where $\Gamma\in\mathbb{R}^{N\times D}$ denotes the measurement matrix, $\xi_{j}\in\mathbb{R}^{D}$ the feature-basis vectors, and $\omega_{j,\mu}\in\mathbb{R}$ the feature weights associated with target $\mu$. When optimization drives the object- and target-specific weights $\omega_{j,\mu}$ to concentrate on the feature patterns of a particular $\mu$, the system exhibits a selective-gain effect: scene measurements become dominated by target-related information flow. These dynamics are expressed as
$\Gamma k\;=\;\begin{bmatrix}\sum_{m=1}^{M}\eta_{1,m}\,\xi_{1}\\[4.0pt] \vdots\\[4.0pt] \sum_{m=1}^{M}\eta_{N,m}\,\xi_{N}\end{bmatrix}\approx\widehat{k}_{\mu},$ (3)
where $\eta_{j,m}$ denotes the contribution of object $m$ to feature channel $j$. Measurements concentrate on the specific target when $\eta_{j,m}\to\delta_{m\mu}$ for each feature channel $j$, i.e., when the contributions from non-target objects vanish.
Functionally, this formulation mirrors the hierarchical transformation from V1 simple cells (edge/texture responses) to IT complex cells (holistic shape responses), with $\Gamma$ playing the role of attentional gain control in V4. V1 encodes local orientations, while V4 integrates global contours under top-down modulation, consistent with our FFC-based frequency-domain focusing. From a functional-analysis perspective, this process can be interpreted as a target-guided modulation on an appropriate Riemannian manifold:
$\Gamma k=\sum_{m=1}^{M}\Gamma\omega_{m}\otimes k_{m}\approx\hat{k}^{\prime}_{0}$ (4)
where $\otimes$ denotes the diffeomorphism operator in the feature space. Establishing a rigorous mathematical model is essential for precise object separation. Consider an input scene containing $N$ latent targets. Sparse representation of individual objects. Each object $k_{i}$ ($i=0,1,\dots,N-1$) can be represented sparsely in a high-dimensional dictionary basis $\Psi$:
$k_{i}=\sum_{j=1}^{M}\psi_{ij}x_{ij}$ (5)
where $\psi_{ij}$ denotes the adaptive sparse basis and $x_{ij}$ the sparse coefficients. Panoramic scene decomposition indicates that the holistic scene can be expressed as a linear combination of objects:
$k=\sum_{i=0}^{N-1}k_{i}=\sum_{i=0}^{N-1}\sum_{j=1}^{M}\psi_{ij}x_{ij}$ (6)
subject to the following constraints: $\|\mathbf{x}_{i}\|_{0}\ll M\quad\text{(Sparsity)},\quad\psi_{ij}\in\mathbb{R}^{D}\quad\text{(Dictionary Basis)},\quad x_{ij}\in\mathbb{R}\quad\text{(Sparse Coefficient)}.$ This representation satisfies the principles of compressed sensing, thereby ensuring object separability in the feature space. This theoretical model is implemented through a bio-inspired neural architecture. The Feedforward Feature Selection Network simulates the ventral (“what”) visual pathway by performing target-sensitive filtering on masked defective inputs $M_{h}$. The Target Reconstruction Network mimics the spatial integration function of the parietal cortex and is dedicated to target reconstruction. The training process is formulated as a physically constrained optimization problem:
$\{\xi,\theta\}=\arg\min\frac{1}{H}\sum_{h=1}^{H}\mathcal{C}(N_{h},\tilde{N}_{h})$ (7)
where $\mathcal{C}(\cdot)$ denotes the loss function. Upon convergence, the FECNet weights $\xi$ encode target knowledge into a feature filter $\phi$. Applying $\phi$ to the panoramic scene yields:
$\phi k=\phi\left(\sum_{i=0}^{N-1}\sum_{j=1}^{M}\psi_{ij}x_{ij}\right)=\sum_{j=1}^{M}\psi_{0j}x_{0j}=\hat{k}_{0}$ (8)
Since $\phi$ is a linear operator, it satisfies the following projection property:
$\phi(x_{ij})=\begin{cases}x_{0j},&\text{if }i=0\\ 0,&\text{if }i\neq 0\end{cases}$ (9)
Assume the dictionary $\Psi$ satisfies the Restricted Isometry Property (RIP), i.e. $\mathrm{RIP}(\Psi)\leq\delta_{2s}$ (separability condition). The underlying mechanism is that the filter $\phi$ implements an (approximate) orthogonal complement in feature space, $\phi=I-P_{\mathrm{bg}}$, thereby suppressing non-target components (i.e., those with $i\neq 0$).
A closed-form estimator for $\phi$ can be obtained by minimizing a least-squares criterion and expressed via the Moore–Penrose pseudoinverse. Defining $k^{(h)}$ as the vectorized scene for sample $h$ and $\widehat{k}_{0}^{(h)}$ as its corresponding target component, we may write
$\phi=\arg\min_{\phi^{\prime}}\sum_{h}\Big\|\hat{k}_{0}^{(h)}-\phi^{\prime}\Big(\sum_{i=0}^{N-1}k_{i}^{(h)}\Big)\Big\|_{F}^{2}$ (10)
$\Rightarrow\quad\phi=\Big(\sum_{h}\hat{k}_{0}^{(h)}\,k^{(h)\,T}\Big)\Big(\sum_{h}k^{(h)}k^{(h)\,T}\Big)^{\dagger}$
where denotes the Moore–Penrose pseudoinverse, and $k^{(h)}\in\mathbb{R}^{D\times 1}$ is the vectorized scene representation for sample $h$. We further constrain $\phi=\mathbf{I}-\mathbf{P}_{\mathrm{bg}}$ as an (approximate) orthogonal projection onto the target subspace.
Equivalently, $\phi$ may be viewed as the best approximation to the ideal projection $I-P_{\mathrm{bg}}$ in a suitable Hilbert space $\mathcal{H}$:
$\phi^{\star}\;=\;\arg\min_{\phi\in\mathcal{H}}\;\big\|\,\phi-(I-P_{\mathrm{bg}})\,\big\|_{F}.$ (11)
Here $\mathcal{H}$ is a Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and induced norm $\|\cdot\|$; $\|\cdot\|_{F}$ denotes the Frobenius norm for matrices, $\|A\|_{F}=\sqrt{\sum_{i}\sum_{j}|a_{ij}|^{2}}$.
Let $\mathbf{q}_{xy}$ and $\mathbf{k}_{ij}$ denote learned query and key feature vectors at spatial locations $(x,y)$ and $(i,j)$, respectively, and let $d$ be the feature dimensionality used for softmax scaling. Under bio-inspired attentional modulation we compute attention weights by
$\alpha_{xy}\;=\;\frac{\exp\!\big(\langle\mathbf{q}_{xy},\mathbf{k}_{ij}\rangle/\sqrt{d}\big)}{\sum_{i,j}\exp\!\big(\langle\mathbf{q}_{xy},\mathbf{k}_{ij}\rangle/\sqrt{d}\big)},$ (12)
where the denominator normalizes over all spatial locations $(i,j)$. When $\mathbf{q}_{xy}$ and $\mathbf{k}_{ij}$ are highly similar, the corresponding attention weight $\alpha_{xy}$ approaches one, consistent with Feature Integration Theory: biased competition in V4 enhances target features via attentional gain.
Require that the Kullback–Leibler divergence between the target and background distributions satisfies
$D_{\mathrm{KL}}\big(P_{\mathrm{target}}\|P_{\mathrm{background}}\big)\;>\;\sum_{i=2}^{M}D_{\mathrm{KL}}\big(P_{\mathrm{target}}\|P_{\mathrm{object}_{i}}\big),$ (13)
where $D_{\mathrm{KL}}(P\|Q)=\sum_{i}P_{i}\log\frac{P_{i}}{Q_{i}}$ denotes the Kullback–Leibler divergence.
By large-deviation principles, when the distributional discrepancy between the target and background dominates inter-object interference, empirical distributions concentrate on distributions minimizing relative entropy; consequently, the estimator tends to concentrate on the target distribution, which implies reduction of the estimated target entropy under suitable conditions. Hence, under these assumptions there exists a unique solution $\widehat{k}_{0}$ such that
$\lim_{t\to\infty}H\big(\widehat{k}_{0}^{(t)}\big)\;=\;H_{\mathrm{target}},$ (14)
where $H(X)=-\sum_{x\in\mathcal{X}}p(x)\log p(x)$ denotes the Shannon entropy.
Under Assumptions A1–A3 (sparsity of object representations, the RIP on dictionary $\Psi$, and bounded measurement noise), there exists a linear operator $\phi$ (learnable by FECNet) that approximates the ideal target projector $I-P_{\mathrm{bg}}$. By compressed-sensing theory, if $\mathrm{RIP}(\Psi)\leq\delta_{2s}$ then the learned projector $\phi$ (approximating $I-P_{\mathrm{bg}}$) admits the error bound
$\|\phi k-\widehat{k}_{0}\|_{2}\;\leq\;C_{1}\,\sigma_{s}(k)\;+\;C_{2}\,\epsilon,$ (15)
where $C_{1}$ and $C_{2}$ depend on the RIP constants and the dictionary condition number; $\sigma_{s}(k)$ denotes the error of the best $s$-term approximation of $k$; and $\epsilon$ denotes an upper bound on the measurement noise.
Together, these results provide a rigorous foundation for the proposed object-centric imaging approach. By casting target separation as a formal projection problem under sparsity constraints and employing information-theoretic concentration arguments to characterize convergence, we show that under the stated assumptions the Specific Object-of-Interest Imaging system provably isolates the target and converges to an optimal reconstruction that up to the approximation and noise terms given above.
3.2 Image Inpainting System
Figure 3: Processing flow of the proposed image-inpainting system.
To enhance structural coherence and reconstruction accuracy for occluded images, we propose a two-stage framework grounded in specific-object-of-interest imaging principles (Fig. 3). Departing from traditional pixel-level completion in RGB space, our architecture adopts an object-first perception inspired by biological vision, decomposing the task into structure-prioritized progressive refinements. Emphasizing semantically coherent reconstruction, we decompose the imaging pipeline into two stages: (1) object perception $\rightarrow$ (2) object inpainting. The core innovation is SIOI, which compels the model to focus on the target object while suppressing irrelevant background interference. This reduces computational overhead by concentrating structural modeling on critical regions (see Fig. 4).
Figure 4: Architecture of the proposed Specific Object-of-Interest Imaging.
Stage I: Specific Object-of-Interest Imaging Biological vision begins with retinal photoreceptors converting incoming light into neural signals. Subsequent processing in the lateral geniculate nucleus (LGN) integrates spatial information and suppresses high-frequency noise via center–surround antagonism. Feature extraction proceeds in primary visual cortex (V1), where simple cells respond to oriented edges and spatial-frequency textures [46]. Object prioritization amplifies target representations through bottom-up attentional biasing [66], suppressing background interference—functionally analogous to selective attention in the ”cocktail-party” effect.
Our SIOI module computationally embodies this hierarchy. We process corrupted inputs with modified U-Net encoders to extract multi-scale features, mimicking V1 orientation selectivity. Embedding Fast Fourier Convolution (FFC) modules in multiple encoding and decoding processes captures long-range dependencies through spectral transformation, simulating attentional modulation in V4 that amplifies target signals. The spectral processing can be written as
$\mathcal{F}_{\mathrm{freq}}\;=\;\mathrm{Conv}_{\mathrm{freq}}\big(\mathrm{FFT2}(\mathcal{F})\big),\qquad\mathcal{F}^{\prime}\;=\;\mathrm{IFFT2}(\mathcal{F}_{\mathrm{freq}}),$ (16)
where $\mathrm{FFT2}$ and $\mathrm{IFFT2}$ denote the 2-D Fourier transform and its inverse, respectively, and $\mathrm{Conv}_{\mathrm{freq}}$ is a learnable convolutional operator applied in the frequency domain.(Fig. 5)
The output of the SIOI module is a preliminary, yet structurally coherent, object map contain local defects. As shown in Fig. 4. Biologically, this parallels V1 edge detection and V4-mediated object-based attention, in which attentional gain enhances target representations while suppressing background activity [74]. Our FFC modules enhance target-relevant spectral components while attenuating background spectra. Functionally, FFC operations correspond to V1-like orientation detection, whereas the learned attention weights emulate pulvinar–thalamic modulation of cortical processing. Consequently, Stage I produces structurally coherent object maps despite occlusions, providing essential priors for subsequent texture synthesis and refinement.This spectral processing overcomes the limitations of standard convolutions for modeling long-range dependencies and significantly improves structural consistency.
Together, these mechanisms reduce information redundancy. The SIOI module serves as a generic, front-end ”visual processor” that transforms a complex, occluded scene into a purified, object-centric representation. It provides a powerful intermediate representation that is the cornerstone of our framework, explicitly containing the structural and semantic priors necessary for high-fidelity inpainting.
Stage II: Object Inpainting (Coarse inpainting + Fine inpainting) To validate the efficacy of the SIOI module independently of advanced inpainting architectures, we deliberately employ a straightforward inpainting network in Stage II. Unlike conventional image inpainting methods, our model eliminates the need for additional manually annotated corruption masks and is capable of automatically identifying and inpainting the corrupted regions of the target. It is critical to emphasize that the innovation of this work is not this network itself, but rather the SIOI module. The purpose of Stage II is solely to demonstrate that the priors generated by SIOI are so potent that they enable even a simple network to achieve high-quality reconstruction. This stage is loosely analogous to the role of higher visual areas in integrating semantic information for detailed perception. Our Structural Recovery Module (SRM) takes the SIOI output and performs contour interpolation and shape completion using multi-scale dilated convolutions and feature-pyramid fusion:
$\mathcal{F}_{\mathrm{MSDC}}\;=\;\mathrm{Concat}\!\left(\bigcup_{r\in\{1,2,4,6\}}\mathrm{Conv}_{r}(x)\right)$ (17)
An FPN-like fusion integrates encoder–decoder hierarchical features, and subpixel-convolution upsampling is used to suppress checkerboard artifacts. The output is a complete structural feature tensor $\mathcal{F}_{\mathrm{SRM}}\in\mathbb{R}^{64\times H\times W}$.
Figure 5: Architecture of the Simple Object Inpainting.
Functionally, both biological and computational systems achieve structural coherence: biological vision generates shape priors via predictive-coding [50], while the SRM fuses multi-scale context through MSDC and feature-pyramid fusion. This design helps to avoid texture blurring and semantic mismatches. Technically, MSDC operations correspond to V2–V3 contour integration, while feature-pyramid fusion approximates cortical feedback modulation. SRM enhances structural integrity in partially occluded object maps from Stage I, enabling a coarse-to-fine imaging strategy.
$\mathcal{F}_{\mathrm{SRM}}\;=\;\mathcal{R}(I_{\mathrm{coarse}}),\qquad\mathcal{F}_{\mathrm{SRM}}\in\mathbb{R}^{64\times H\times W},$ (18)
where $\mathcal{R}(\cdot)$ denotes SRM processing and $I_{\mathrm{coarse}}$ denotes the coarsely restored map. Subsequently, the Global Detail Recovery Module (GDRM) utilizes a symmetric encoder-decoder with skip connections to synthesize fine-grained textures onto the recovered structure, guided by the robust priors from Stage I. Our computational counterpart, the Global Detail Recovery Module, implements a symmetric encoder–decoder architecture that consumes the 64-channel structural features $\mathcal{F}_{\mathrm{SRM}}\in\mathbb{R}^{64\times H\times W}$ produced by the SRM. This simple design choice underscores our core claim: The SIOI module is a versatile, plug-and-play component that can elevate the performance of any existing inpainting pipeline by providing superior object-specific guidance.
3.3 Loss function
To achieve high-fidelity reconstruction of defective object images, we design a multi-branch fused loss that jointly promotes structural coherence, fine-detail reconstruction, and global visual-style consistency. The composite loss comprises three components: a weighted reconstruction loss [48, 75], a perceptual loss [76], and a style loss [77]. Integrated with principled weighting and multi-stage optimization strategies, the loss ensures that synthesized images exhibit improved semantic structure and textural fidelity.
Weighted Reconstruction Loss. We introduce a spatially weighted reconstruction loss based on target maps. An explicit weight map $w(x,y)$ differentially weights per-pixel reconstruction errors in the $L_{1}$$L_{2}$ terms:
$\mathcal{L}_{\mathrm{weighted}}\;=\;\frac{1}{N}\sum_{x,y}\big(\alpha\,w(x,y)+1\big)\,\big\|\hat{I}(x,y)-I_{\mathrm{gt}}(x,y)\big\|_{2}^{2}$ (19)
Equivalently, at pixel-level the reconstruction loss can be written as
$\mathcal{L}_{\mathrm{recon}}\;=\;\sum_{i}w_{i}\,\big\|I_{i}^{\mathrm{pred}}-I_{i}^{\mathrm{gt}}\big\|_{2}^{2},$ (20)
where $\hat{I}(x,y)$ denotes the predicted intensity at pixel $(x,y)$, $I_{\mathrm{gt}}(x,y)$ the corresponding ground truth, $I_{i}^{\mathrm{pred}}$ and $I_{i}^{\mathrm{gt}}$ the RGB vectors at pixel $i$, $w_{i}$ the pixel-level weight, and $\alpha$ a scalar weighting factor (empirically set to 4). Higher weights are assigned to semantic target regions and lower weights to background regions; this directs the network to prioritize detail reconstruction in structurally critical areas and effectively mitigates edge blurring and structural discontinuities. This mechanism intensifies supervision in target regions while reducing gradient interference from irrelevant areas.
Perceptual Loss. To compensate for pixel-level losses’ inability to capture high-level semantics, we incorporate a perceptual loss computed from feature activations of a pretrained VGG-16 network:
$\mathcal{L}_{\mathrm{perc}}\;=\;\sum_{l\in\mathcal{L}}\frac{1}{C_{l}H_{l}W_{l}}\,\big\|\phi_{l}(I^{\mathrm{pred}})-\phi_{l}(I^{\mathrm{gt}})\big\|_{2}^{2},$ (21)
where $\phi_{l}(\cdot)$ denotes the feature maps extracted from the $l$-th VGG-16 layer, and $C_{l},H_{l},W_{l}$ denote the channel count, height, and width of $\phi_{l}$, respectively. By comparing representations across selected VGG-16 layers, the perceptual loss captures high-level semantics (e.g., object shapes and textures) and enhances semantic coherence in generated objects.
Style Loss. The style loss quantifies texture distribution and stylistic characteristics using Gram matrices as in style-transfer methods [77]. For feature maps $\phi_{l}(x)$ at layer $l$ of VGG-16, the Gram matrix is
$G_{l}(x)\;=\;\phi_{l}(x)\,\phi_{l}(x)^{\top},$ (22)
and the style loss is
$\mathcal{L}_{\mathrm{style}}\;=\;\sum_{l\in\mathcal{L}}\big\|G_{l}(I^{\mathrm{pred}})-G_{l}(I^{\mathrm{gt}})\big\|_{F}^{2},$ (23)
where $\mathcal{L}$ denotes the set of selected layers and $\|\cdot\|_{F}$ the Frobenius norm. This statistical matching captures local texture statistics and global stylistic discrepancies, improving the naturalness of restored images and reducing pattern repetition and artifacts.
Composite Loss. The integrated loss combines the above terms. The weighted reconstruction term enhances edge clarity and structural continuity in semantic regions; the perceptual term reinforces high-level semantic consistency aligned with human perception; and the style term mitigates texture artifacts across diverse objects and scenes:
$\mathcal{L}_{\mathrm{total}}=\lambda_{1}\mathcal{L}_{\mathrm{recon}}+\lambda_{2}\mathcal{L}_{\mathrm{perc}}+\lambda_{3}\mathcal{L}_{\mathrm{style}}\quad(\lambda_{1}=1.0,\,\lambda_{2}=0.1,\,\lambda_{3}=250)$ (24)
These weights are empirically chosen to balance the differing numerical scales and competing objectives of the loss components during training.
4 Experiment
4.1 Dataset
Four distinct datasets were employed. A custom object-centric teapot dataset (2,000 training, and 200 test samples) constructed in our laboratory featuring $512\times 512$ scene images of multi-view teapot objects against complex backgrounds with artificial items and indoor environments, each paired with corresponding $512\times 512$ ground-truth masks that underwent multi-stage manual calibration to preserve geometric structures and textural details; alongside three COCO2017-derived subsets—“elephant” (2,143 training images, 89 validation as test), “giraffe” (2,546 training, 101 validation as test), and “zebra” (1,916 training, 85 validation as test)—all uniformly resized to $512\times 512$ resolution to ensure comparability. A consistent target-aware occlusion protocol was implemented across all datasets, automatically generating masks from target maps to restrict occlusions exclusively to foreground areas, with coverage constrained to $\sim$40% of object regions during both training and inference phases to preserve contextual information while necessitating complex structural reasoning.
4.2 Experimental Setup
Implementation Details. All experiments were conducted on an Ubuntu 20.04 system equipped with an Intel Xeon Platinum 8358P CPU and an NVIDIA GeForce RTX 4090 GPU (24 GB VRAM). The framework was implemented in PyTorch 1.10.0 with CUDA 11.3 support. Training spanned 50 epochs with a batch size of 4, constrained by available computational resources. Optimization employed Adam with a learning rate of $2\times 10^{-4}$, resulting in approximately 0.5 hours of total training time on the specified hardware. We adopted a two-stage training protocol: First, the ObjectImagingModule was pretrained exclusively using $\mathcal{L}_{\mathrm{recon}}$ to generate semantically structured intermediate representations. This plug-and-play module was subsequently frozen and integrated across all four experimental configurations. Second, the Structural Refinement Module (SRM) and Geometric Detail Reconstruction Module (GDRM) were trained using $\mathcal{L}_{\mathrm{perc}}$ and $\mathcal{L}_{\mathrm{style}}$ to enhance semantic coherence and stylistic fidelity. This secondary optimization (implemented solely for preliminary validation in Experiment 1) maintained the ObjectImagingModule’s parameters fixed. Full implementation code is available at: https://github.com/WYH302/WuYonghao.git.
Baseline. To rigorously evaluate the efficacy of our proposed method, we benchmark against a representative spectrum of state-of-the-art inpainting approaches selected through a multi-faceted lens. Our criteria prioritize comprehensive coverage of methodological evolution, spanning from classical exemplar-based techniques to cutting-edge generative models, while ensuring architectural diversity across diffusion, transformer, and GAN-based paradigms. The evaluation suite begins with foundational exemplar-based methods: Criminisi [18], representing classical non-local patch synthesis, establishes a performance baseline for traditional non-deep-learning approaches and highlights inherent limitations in complex structural recovery. Its optimized counterpart, PatchMatch [78], provides critical insights into computational efficiency trade-offs through accelerated nearest-neighbor matching. Transitioning to modern deep learning, AOT-GAN [79] is included for its attention-driven, long-range dependency modeling via dilated convolutions—directly addressing holistic scene coherence challenges central to our work. Edge-Connect [56] offers a strategic comparative lens as a two-stage edge-guided model, enabling explicit evaluation of structural prior utilization against our coarse-to-fine refinement paradigm. The hybrid methodology of MISF [80] tests multi-scale feature interaction between traditional and deep components against our biologically inspired framework. Finally, contemporary diffusion models LaMa [81] and RePaint [41] are incorporated as cutting-edge references, with LaMa validating spectral processing choices and RePaint contrasting stochastic refinement against our deterministic pipeline. All baselines underwent five independent runs with randomized initializations, utilizing official implementations and pre-trained models where available.
Evaluation metrics. We employed four widely used metrics to quantify inpainting performance: (1) SSIM (Structural Similarity Index Measure), which assesses structural similarity between predictions and ground truth (higher is better); (2) PSNR (Peak Signal-to-Noise Ratio), measuring pixel-wise reconstruction fidelity (higher is better); (3) MAE (Mean Absolute Error), which measures average absolute pixel deviation (lower is better); and (4) LPIPS [82], a learned perceptual metric based on deep features (lower is better).
4.3 Preliminary Theoretical Verification
To isolate and validate the core contribution of our work—the Specific Object-of-Interest Imaging (SIOI) module—we first conducted a controlled experiment using a simplistic subsequent inpainting network (Stage II: SRM+GDRM). It is crucial to emphasize that this inpainting network is intentionally simplistic and is not the focus of our innovation. Its sole purpose is to serve as a stable and transparent baseline to demonstrate that the high-quality output is primarily attributable to the superior structural and semantic priors generated by the SIOI module (Stage I).
Unlike conventional image inpainting methods, our model eliminates the need for additional manually annotated corruption masks and is capable of automatically identifying and inpainting the corrupted regions of the target. We deliberately retained some black borders to demonstrate that the model does not solely rely on the color black to identify corrupted regions. As shown in Table 2 and Fig. 6, even with this simple network, the pipeline achieves exceptional performance on the Teapot dataset. This result provides foundational evidence that the SIOI module successfully accomplishes its design goal: to extract a clean, well-defined, and structurally sound representation of the target object from a complex and occluded scene. This output serves as an ideal input for anydownstream inpainting task, thereby validating the SIOI module’s efficacy as a standalone component.
Table 2: Preliminary Theoretical Verification on the teapot dataset (mean $\pm$ std over five runs). ↑ means higher is better, and ↓ means lower is better.
[…]
Experimental values $0.98\pm 0.02$ $33.86\pm 0.38$ $1.61\pm 0.18$ $0.02\pm 0.01$
Figure 6: Preliminary Theoretical Verification of exemplar inpainting results (teapot dataset).
4.4 Generalized empirical evaluation——Quantitative comparisons
Having established the standalone effectiveness of the SIOI module, we now evaluate its primary value proposition: its utility as a universal, plug-and-play front-end that can enhance diverse, off-the-shelf inpainting models. We integrated our SIOI module with a wide spectrum of state-of-the-art methods, spanning from classical algorithms to modern deep learning approaches. To ensure a fair comparison, all baseline methods were re-evaluated using a consistent object-level evaluation protocol, where their outputs were multiplied by corresponding object masks prior to quantitative analysis.
As summarized in Table 3, the SIOI module consistently enables significant performance improvements across all evaluation metrics. The key observation is not the absolute performance of any single model, but the transformative gain achieved by prepending SIOI. For instance, SIOI enables Criminisi to achieve a PSNR of 31.552 dB, a 7 dB improvement over the vanilla version. More notably, SIOI boosts LaMa to a staggering 45.222 dB PSNR and 0.0023 LPIPS, significantly surpassing all other methods. The improvements are especially pronounced on the Teapot dataset, where SIOI+LaMa achieves near-perfect reconstruction (SSIM: 0.9973, LPIPS: 0.0023).These results demonstrate that SIOI is not merely complementary but revolutionary, allowing both traditional and modern algorithms to transcend their inherent limitations by providing a purified, object-centric input. Regarding computational efficiency, SIOI acts as a generic adapter, effectively bridging the gap between holistic image processing and efficient object-centric reconstruction.
Across multiple datasets, models incorporating SIOI consistently achieve superior results in SSIM, PSNR, MAE, and LPIPS metrics compared to those without SIOI. Furthermore, models employing SIOI exhibit significantly enhanced perceptual quality. These results substantiate the effectiveness of SIOI in achieving high-fidelity reconstruction of objects with complex textures. Additionally, the object-centric imaging paradigm facilitates consistent separation of target objects across diverse backgrounds. Regarding computational efficiency, SIOI reduces redundant computations by concentrating processing resources on target regions, thereby accelerating inference while maintaining reconstruction quality.
Table 3: Experimental Results for Different Datasets and Models, ↑ means higher is better, and ↓ means lower is better.
Datasets Model SSIM $\uparrow$ PSNR $\uparrow$ MAE $\downarrow$ LPIPS $\downarrow$
Teapot Criminisi 0.9415 24.597 3.5517 0.0965
PatchMatch 0.9212 21.219 5.4619 0.1263
Edge-Connect 0.9457 25.095 3.3687 0.0908
AOT-GAN 0.9487 25.329 3.2564 0.0863
MISF 0.9370 24.871 3.7130 0.1178
LaMa 0.9524 25.534 3.1024 0.0833
RePaint 0.9330 24.037 4.1087 0.1141
SIOI+Criminisi 0.9841 31.552 0.7396 0.0246
SIOI+PatchMatch 0.9752 25.912 1.7894 0.0408
SIOI+Edge-Connect 0.9795 32.378 0.8364 0.0299
SIOI+AOT-GAN 0.9248 36.940 1.0471 0.0094
SIOI+MISF 0.9739 33.487 1.0541 0.0674
SIOI+LaMa 0.9973 45.222 0.1119 0.0023
SIOI+RePaint 0.9712 28.599 1.6907 0.0645
elephant Criminisi 0.9671 32.615 1.3453 0.0262
PatchMatch 0.9596 28.643 2.1005 0.0441
Edge-Connect 0.9562 31.958 1.7455 0.0423
AOT-GAN 0.8416 27.653 4.9353 0.0608
MISF 0.8472 23.789 5.2913 0.1908
LaMa 0.9236 26.059 2.9596 0.0949
RePaint 0.8624 23.690 5.1769 0.1676
SIOI+Criminisi 0.9774 33.152 0.8312 0.0190
SIOI+PatchMatch 0.9721 29.327 1.3736 0.0325
SIOI+Edge-Connect 0.9636 30.693 1.2479 0.0448
SIOI+AOT-GAN 0.9341 34.691 1.4350 0.0243
SIOI+MISF 0.8883 29.070 3.4237 0.1848
SIOI+LaMa 0.9786 34.706 0.6523 0.0140
SIOI+RePaint 0.9102 28.956 3.2348 0.1489
giraffe Criminisi 0.9721 29.962 1.5050 0.0184
PatchMatch 0.9653 27.671 2.3031 0.0368
Edge-Connect 0.9678 29.914 1.6754 0.0280
AOT-GAN 0.9109 26.742 4.1489 0.0388
MISF 0.8998 23.505 4.6733 0.1208
LaMa 0.9424 25.732 2.5901 0.0670
RePaint 0.9079 23.563 4.5075 0.1096
SIOI+Criminisi 0.9788 29.824 1.1000 0.0170
SIOI+PatchMatch 0.9734 26.327 1.7493 0.0306
SIOI+Edge-Connect 0.9676 27.527 1.6295 0.0382
SIOI+AOT-GAN 0.8741 30.861 2.3573 0.0350
SIOI+MISF 0.9325 27.270 3.0461 0.1215
SIOI+LaMa 0.9796 31.023 0.9491 0.0129
SIOI+RePaint 0.9465 27.109 2.8830 0.0992
zebra Criminisi 0.9677 27.598 2.2333 0.0236
PatchMatch 0.9262 22.627 4.1497 0.0857
Edge-Connect 0.9588 27.891 3.1040 0.0263
AOT-GAN 0.8451 22.399 9.6206 0.0486
MISF 0.8583 21.409 8.2618 0.1596
LaMa 0.9331 24.284 3.7468 0.0684
RePaint 0.8677 21.544 8.0205 0.1499
SIOI+Criminisi 0.9750 27.389 1.6940 0.0211
SIOI+PatchMatch 0.9675 24.691 2.4553 0.0365
SIOI+Edge-Connect 0.9632 26.089 2.2360 0.0387
SIOI+AOT-GAN 0.9070 28.025 3.2725 0.0527
SIOI+MISF 0.8960 24.048 6.2437 0.1556
SIOI+LaMa 0.9772 28.811 1.3896 0.0144
SIOI+RePaint 0.9130 24.309 5.9439 0.1365
4.5 Generalized empirical evaluation——Qualitative comparisons
Qualitative results further corroborate the role of SIOI as a powerful pre-processing stage. By providing a clean, background-suppressed, and structurally coherent representation of the target object, SIOI enables even conventional algorithms to produce more semantically consistent and visually plausible results. Fig. 7, Fig. 8, Fig. 9, Fig. 10, Fig. 11, Fig. 12, Fig. 13, Fig. 14, presents qualitative results that highlight characteristic behaviors and failure modes of the benchmark models. Criminisi [18] fundamentally lacks semantic understanding in its non-local patch matching, leading to severe structural discontinuities and semantically inconsistent textures in complex scenes. Similarly, PatchMatch [78] remains constrained by hand-crafted features and low-level metrics, failing to capture high-level semantics or global coherence, which results in visible patch artifacts—a byproduct of its localized/static patch-level processing that becomes evident under gradient-rich lighting and on non-rigid surfaces. AOT-GAN [79] exhibits jagged artifacts and discontinuous textures in complex regions, these failures stem from its use of dilated convolutions for feature aggregation. Edge-Connect [56], while effective in many cases, suffers from visible spatial misalignment due to its over-reliance on imperfect Canny edges, which do not correspond to perceptually relevant structures and disrupt spatial coherence in repetitive textures (e.g., grass). MISF [80] produces overly smooth, artificial-looking edges and illumination inconsistencies, resulting from its multi-scale regularization and strong smoothing priors as well as incompatible traditional and deep feature interactions. LaMa [81] over-smooths detailed regions and produces blurry or insufficiently sharp outputs despite its efficient Fourier-based design. RePaint [41] often yields overly smooth or blurry outputs with incongruous textures due to its slow. Without using our method, existing methods exhibit characteristic failures due to background interference and a lack of object-level understanding, resulting in color and texture details related to the background that are completely inconsistent with their original objects.
In contrast, the integration of the SIOI module consistently guides the inpainting focus, resulting in improved edge sharpness, texture consistency, and semantic coherence, regardless of the backend inpainting algorithm used. Experimental results across multiple object datasets show that our approach performs particularly well on rigid objects (e.g., teapots), while achieving relatively weaker results on non-rigid objects such as giraffes, zebras, and elephants. We attribute this difference to the higher morphological variability and structural complexity of non-rigid categories, which may require larger or more adaptive model parameters to adequately capture their semantic and textural characteristics. Preliminary observations suggest a potential linear relationship among semantic complexity, training data scale, and model parameter quantity—a phenomenon that merits further quantitative investigation in future work.
Figure 7: Qualitative comparison of exemplar inpainting results (teapot dataset).
Figure 8: Qualitative comparison of exemplar SIOI inpainting results (teapot dataset).
Figure 9: Qualitative comparison of exemplar inpainting results (elephant dataset).
Figure 10: Qualitative comparison of exemplar SIOI inpainting results (elephant dataset).
Figure 11: Qualitative comparison of exemplar inpainting results (giraffe dataset).
Figure 12: Qualitative comparison of exemplar SIOI inpainting results (giraffe dataset).
Figure 13: Qualitative comparison of exemplar inpainting results (zebra dataset).
Figure 14: Qualitative comparison of exemplar SIOI inpainting results (zebra dataset).
4.6 Ablation Studies
In fact, a detailed ablation experiment has already been conducted above, So this is just a simple supplement to the Preliminary Theoretical Verification experiment, also conducted on the Teapot dataset. This experiment is based on the preliminary theoretical verification experiment. To evaluate the effectiveness of our methodology, we conducted ablation studies under identical experimental conditions. Key components were systematically removed, and the resulting imaging outcomes were comparatively analyzed to elucidate performance differences. As shown in Fig. 15.
Figure 15: Illustration of ablation experiment.
Ablation I: In Ablation I, the state I network was removed while retaining the object-centric input–output pipeline. The corresponding results are presented in Tab. 4. The resulting model showed marginal superiority in certain quantitative metrics compared to the proposed method—SSIM (0.976 vs. 0.978), PSNR (33.95 dB vs. 33.86 dB), MAE (1.313 vs. 1.605), and LPIPS (0.017 vs. 0.018)—but these differences were statistically insignificant in terms of perceptual quality. However, subjective evaluation revealed compromised edge localization, manifested as jagged artifacts and global blurring. This degradation confirms the critical role of Stage I in achieving high structural precision. Furthermore, visual comparisons clearly demonstrated the superior inpainting quality achieved when Stage I was integrated. The marginal improvements observed in PSNR and MAE, despite perceptual and qualitative degradation, further highlight that such metrics can be misleading indicators of inpainting quality, as they from the loss function’s direct penalization of RGB deviations, do not consistently align with human perceptual judgment.
Ablation II: In Ablation II, the object-centric I/O was replaced with full-scene inputs. The model processed the entire image, including the extensive and complex background, which resulted in severe performance degradation. This represents the standard paradigm used by most existing methods, which our SIOI approach aims to supersede for object-centric tasks. Quantitative metrics showed substantial deterioration: SSIM: 0.964, PSNR: 29.43 dB, MAE: 2.085, LPIPS: 0.065. Visual inspection confirmed that this configuration yielded the poorest inpainting quality among all compared models. It exhibited pronounced detail loss, extensive artifacts throughout the restored regions, and markedly impaired image clarity. Furthermore, the model’s inability to effectively extract features, particularly from the edges of degraded regions, resulted in insufficient feature representation during inpainting. This shortfall directly led to blurry reconstruction and pervasive artifact generation. This ablation highlights the fundamental advantage of our object-first paradigm: it dramatically reduces the solution space by focusing computation exclusively on the target, thereby avoiding the pitfalls of whole-image processing.
In contrast to both ablated versions, the full model (with SIOI) achieved superior performance, delivering sharp edges, semantically coherent structures, and photorealistic textures. This provides compelling evidence that our two-stage pipeline, headed by the SIOI module, effectively addresses the bottlenecks of traditional inpainting by first resolving the ”where and what” of the target object before addressing the ”how” of its inpainting. The SIOI module acts as a powerful pre-processor that can be plugged into any inpainting network to elevate its performance on object-centric tasks.
Table 4: Comparison of ablation study results on the Teapot dataset. (mean $\pm$ std over five runs). ↑ means higher is better, and ↓ means lower is better.
[…]
Origin $0.98\pm 0.01$ $33.86\pm 0.38$ $1.61\pm 0.08$ $0.02\pm 0.01$
Ablation I $0.98\pm 0.01$ $33.95\pm 0.19$ $1.31\pm 0.06$ $0.02\pm 0.01$
Ablation II $0.96\pm 0.01$ $29.43\pm 0.25$ $2.09\pm 0.15$ $0.07\pm 0.02$
4.7 Robustness Analysis
This experiment is based on the preliminary theoretical verification experiment, also conducted on the Teapot dataset. To evaluate the generalization ability of our method under challenging conditions, we conducted robustness experiments on the following scenarios:
  • 1.
    Low illumination: 50% reduced scene brightness
  • 2.
    Small targets: Objects occupying $5\%$ image area
  • 3.
    Gaussian noise: Additive noise ($\mu=0,\sigma=1$)
  • 4.
    Salt-and-pepper noise: Impulse noise ($p_{\text{salt}}=0.01,p_{\text{pepper}}=0.01$)
  • 5.
    Multi-object occlusion: 3 instances
  • 6.
    Motion blur: Linear kernel with $15\times 15$ pixels
It should be noted that these challenging scenarios were not part of the training data. For qualitative assessment, two representative samples were selected per category. Quantitative evaluation was omitted due to the limited number of available samples for each specific condition, as illustrated in Fig. 16.
The qualitative results demonstrate the SIOI module’s notable robustness in several challenging conditions. It consistently maintained satisfactory target extraction performance under low illumination, small targets, salt-and-pepper noise, and motion blur. This resilience can be attributed to the module’s design principles: its spectral processing (FFC) is inherently robust to global illumination changes and motion artifacts, while its attentional mechanisms, inspired by V4 cortical processing, enable it to amplify and track small or salient targets despite noise. Performance degradation emerged in scenarios with three or more coexisting objects, as the competition for attentional resources increased. This observation directly mirrors the biological phenomenon of attentional limitation described in Desimone and Duncan’s biased competition model [66], validating the bio-inspired nature of our approach even in its failure modes.
The most substantial degradation occurred under high-power Gaussian noise. This is a recognized challenge for feature-based attentional systems both in biological and artificial networks. The SIOI module’s reliance on gradient and frequency information for target selection makes it particularly susceptible to this type of non-structured, high-entropy noise, which overwhelms the salient features the module is designed to detect. This result precisely highlights the critical role and inherent vulnerability of the front-end SIOI module: its output quality is the foundational prior for all subsequent processing. This underscores the importance of our modular design—the failure is contained and diagnosed at the SIOI stage, preventing a cascading collapse of the entire system and clearly pointing to the component that requires improvement.
These findings provide a clear roadmap for evolving the SIOI module. Future work will focus on enhancing its robustness against unstructured noise, potentially by incorporating neuromorphic noise-adaptation mechanisms inspired by the auditory brainstem or retinal adaptive gain control. This will further solidify the SIOI module’s role as a robust, plug-and-play pre-processor for object-centric vision tasks.
In summary, the robustness analysis confirms that the SIOI module provides a reliable and highly effective object-specific prior across a wide spectrum of challenging conditions. Its performance profile—strong in structured degradation but susceptible to unstructured noise—precisely characterizes its operational envelope and offers valuable insights for its application and further development as a general-purpose biological vision-inspired front end.
Figure 16: Illustration of robustness experiment.
5 Discussion
Image impainting is a job without a standard answer and often leads to controversy, as shown in Fig. 17. We don’t even know if the black area should be an buffalo head or a zebra head. The same model, when the same image is input multiple times, will yield different results. And the work proposed in this article can solve this problem. This work introduces Specific Object-of-Interest Imaging (SIOI)—a novel, bio-inspired paradigm that fundamentally rethinks the pipeline for object-centric visual computing tasks. Departing from conventional holistic processing, SIOI provides a plug-and-play front-end module that first isolates and enhances the target object, generating a refined structural and semantic prior that can be seamlessly integrated into any subsequent processing network. Our experimental validation demonstrates that incorporating the SIOI module consistently elevates the performance of diverse inpainting architectures. This is not a contest between inpainting models; rather, it is a demonstration that providing a high-quality, object-centric prior is a more impactful lever for improvement than designing yet another holistic inpainting network.
Figure 17: Misunderstand of image impainting.
This approach aligns with the ventral pathway’s ”object-before-background” processing strategy, where attentional mechanisms in areas like V4 selectively amplify target features while suppressing irrelevant background clutter[83]. Our SIOI module computationally instantiates this principle, thereby mitigating the fundamental issue of information redundancy that plagues whole-image methods. Compared to GAN- or diffusion-based approaches that operate on the entire scene, our method avoids propagating errors from background regions by focusing computation on the target, akin to V4’s role in contour integration and feature binding.
The robustness of the SIOI module under extreme conditions (e.g., low illumination, noise) further underscores its utility as a reliable pre-processor. Its behavior is consistent with the object-based attention mechanism in biological vision, which dynamically modulates early visual areas (V1-V4) to maintain target salience amidst distraction. It is important to note that the performance of any subsequent model is contingent on the quality of the prior provided by SIOI. This dependency mirrors the dorsal-ventral synergy in the brain, where spatial localization guides object recognition. This highlights that advancing object-level pre-processing is a critical frontier in itself.
We emphasize that the inpainting network presented in Stage II serves primarily as a proof-of-concept vehicle to validate the efficacy of the SIOI prior. our architecture may not represent the ultimate performance ceiling, and that future researchers may develop superior networks. The paramount innovation of this work is not this specific network, but the SIOI paradigm itself. This paradigm shift—from direct holistic inpainting to object-first prior generation—is our central contribution. The generalization of the SIOI module is inherently tied to the semantic definition of a ”specific object.” This is not merely a limitation but a defining characteristic and a research opportunity. But we must frankly admit that the generalization of this technology is not strong due to its focus on specific objects. And each semantic object represents a type of dataset, which also requires retraining the object imaging network. For example, whether cat is an object, or is it divided into jungle cat felis chaus, sand cat felis Margarita, black foot cat felis nigripes, and wild cat felis silvestris(which is also more refined as a specific object) needs further research. The module’s performance can be scaled and specialized, offering a practical pathway for technology transfer, allowing the SIOI paradigm to be precisely tailored to application-specific needs.
Future work will explore integrating prefrontal-cortex-like task-driven modulation to enable dynamic reconfiguration of the SIOI module for different object categories and end goals, advancing toward truly adaptive and biologically plausible computational vision systems.
6 Conclusion
In conclusion, this work proposes Specific Object-of-Interest Imaging (SIOI) as a novel, bio-inspired pre-processing paradigm for object-centric visual tasks. SIOI operates by first isolating and enhancing the target object, providing a high-fidelity prior that dramatically improves the performance of subsequent vision modules. Extensive experiments confirm that the SIOI module is a powerful, plug-and-play front-end that consistently improves various inpainting models by suppressing background interference and providing a coherent object-centric prior. Its design is grounded in the ”object-first” processing hierarchy of the ventral visual pathway, offering a computational solution to the problem of information redundancy.The framework demonstrates strong robustness across challenging conditions, reflecting the resilience of biological attentional systems. Limitations arise from reliance on initial object generation, which may impair boundary guidance. We envision the SIOI paradigm extending beyond inpainting to benefit a wide array of object-centric applications, including segmentation, detection, and video tracking, wherever suppressing background clutter and focusing computational resources on the target is paramount. Future work will explore task-driven modulation and fMRI validation, advancing toward biologically adaptive imaging and broader applications in neuroimaging and perceptual assistance.
Table 5: Extended Mathematical Notation Guide with Biological Correlations
Symbol Mathematical Definition Biological Mechanism Network Implementation
$\Gamma$ Target-sensitive measurement matrix: $\Gamma k\approx\sum_{m=1}^{M}(\Gamma\omega_{m})\otimes k_{m}$ LGN attentional filtering SIOI spectral filtering
$k$ Scene tensor: $k\in\mathbb{R}^{H\times W\times C}$ Retinal photoreceptor output encoding SIOI degraded input (Stage I)
$\otimes$ Diffeomorphic operator: $(\Gamma\omega_{m})\otimes k_{m}$ Foveal cortical magnification FFCModule feature warping
$\xi_{j}$ Orthogonal feature basis: $\xi_{j}\in\mathbb{R}^{D}$ V1 simple cell orientation responses Multi-scale encoder features in SIOI
$\omega_{m}$ Feature weights: $\omega_{m}\rightarrow\delta_{m\mu}$ Neural resource allocation Attention weights
$k_{i}$ Latent object tensor: $k_{i}\in\mathbb{R}^{h_{i}\times w_{i}\times C}$ IT cortex semantic object representation Target object map (Stage I output)
$\phi$ Projection operator: $\phi=I-P_{\mathrm{bg}}$ Dorsal pathway spatial suppression TRNet background suppression
$D_{KL}$ Kullback–Leibler divergence: $D_{KL}(P\|Q)=\sum_{i}P_{i}\log\frac{P_{i}}{Q_{i}}$ Neural coding efficiency optimization Loss function convergence
$\mathcal{F}$ Spectral feature map: $\mathcal{F}^{\prime}=\mathrm{IFFT2}\!\left(\mathrm{Conv_{freq}}\!\left(\mathrm{FFT2}(\mathcal{F})\right)\right)$ V1–V4 spatial frequency processing FFCModule operations in bottleneck
$\mathcal{L}_{\mathrm{total}}$ Composite loss: $\mathcal{L}_{\mathrm{total}}=\lambda_{1}\mathcal{L}_{\mathrm{recon}}+\lambda_{2}\mathcal{L}_{\mathrm{perc}}+\lambda_{3}\mathcal{L}_{\mathrm{style}}$ Prefrontal cortex task-driven modulation Phased training strategy

### v3

@@
-Image Inpainting based on Visual-Neural-Inspired Specific Object-of-Interest Imaging Technology
+Visual-Neural-Inspired Image Inpainting for Specific Objects-of-Interest Imaging
 Abstract
-Conventional image inpainting methods operate on whole images.
-Drawing on cortical processing principles this study addresses the bottlenecks of conventional holistic image inpainting methods–susceptibility to informational redundancy and low computational efficiency under occlusions and complex backgrounds, by proposing a novel framework: ”Specific Object-of-Interest Imaging”.
-This stage extracts and encodes object-level representations from complex scenes, generating semantic and structural priors that can be seamlessly integrated into any inpainting framework.
-On our Teapot object dataset, Elephant object dataset, Giraffe object dataset and Zebra object dataset experimental validation demonstrates–compared with the repair model without using this method, the repair model using this method has more advantages, across metrics including SSIM, PSNR, MAE, and LPIPS, while maintaining robustness in extreme scenarios (low illumination, high noise, multi-object occlusion, motion blur).
-Theoretical analysis integrated with cognitive neuroscience perspectives reveals profound correlations between the ”object precedence perception” mechanism and dynamic feature modulation in visual cortices (V1–V4).
-We demonstrate that incorporating Stage I significantly enhances object consistency and semantic coherence across various mainstream inpainting models.
-This work not only achieves efficient and precise target-centric imaging but also pioneers interdisciplinary pathways bridging brain-inspired computational frameworks with advanced image inpainting techniques.
+Conventional image inpainting techniques typically process entire images, which often leads to computational inefficiency and susceptibility to information redundancy, particularly in occluded or cluttered scenes.
+Inspired by cortical processing mechanisms, this study introduces a novel framework termed “Specific Object-of-Interest Imaging” (SIOI) to overcome these limitations.
+The proposed approach first extracts and encodes object-level representations from complex scenes, producing structural and semantic priors that can be seamlessly integrated into existing inpainting pipelines.
+Extensive evaluations on dedicated object datasets—Teapot, Elephant, Giraffe, and Zebra—demonstrate that models incorporating SIOI consistently outperform those without it across key metrics including SSIM, PSNR, MAE, and LPIPS.
+The framework also exhibits strong robustness under challenging conditions such as low illumination, high noise, multiple object occlusions, and motion blur.
+Furthermore, theoretical analysis grounded in cognitive neuroscience reveals meaningful connections between the “object-first” perceptual mechanism and dynamic feature modulation in visual cortical areas (V1–V4).
+By significantly enhancing object consistency and semantic coherence across diverse inpainting models, this work not only advances efficient and target-aware image restoration but also fosters interdisciplinary convergence between brain-inspired computation and cutting-edge inpainting technologies.
 keywords:
 Image Inpainting , Specific Object-of-Interest Imaging , Visual Neural Mechanisms , Deep Neural Network
-journal: Neural Networks
 orrefcempty
 […]
-organization=School of Electronic Science and Engineering (School of Microelectronics), South China Normal University, addressline=Foshan, 528000, country=China
+[1]organization=School of Electronic Science and Engineering (School of Microelectronics), South China Normal University, addressline=Foshan, 528000, country=China
 […]
-organization=St. Petersburg School of Shipbuilding and Ocean Technology, Guangdong Ocean University, addressline=Zhanjiang 524088, country=China
+[2]organization=St. Petersburg School of Shipbuilding and Ocean Technology, Guangdong Ocean University, addressline=Zhanjiang 524088, country=China
 […]
-organization=Robotics Laboratory, Institute of Automation and Control Processes, addressline=Far Eastern Branch of the Russian Academy of Sciences, city=Vladivostok, postcode=690041, country=Russia
-[…]
-organization=Robotic Laboratory, Institute of Automation and Control Processes FEB RAS, city=Vladivostok, postcode=690041, country=Russia
+[3]organization=Robotics Laboratory, Institute of Automation and Control Processes, addressline=Far Eastern Branch of the Russian Academy of Sciences, city=Vladivostok, postcode=690041, country=Russia
 {graphicalabstract}
@@
 1 Introduction
-Image Inpainting, a foundational and critically important task in computer vision and image processing, aims to restore missing or corrupted regions.
-Such impairments may arise from accidental occlusions during capture, sensor-induced noise, data loss in encoding/decoding or transmission, intentional object removal in post-processing, or physical deterioration in historical photographs.
-The capacity to restore image integrity, visual authenticity, and semantic plausibility renders this technology indispensable across diverse applications, including cultural heritage preservation, digital photography enhancement, medical image reconstruction, cinematic visual effects production, and the interpretation of occluded scenarios in autonomous driving systems.
-However, producing highly accurate and semantically coherent inpainting results remains a fundamental challenge.
-Conventional inpainting methods primarily rely on intrinsic statistical properties and local structural information of images [1, 2, 3, 4, 5, 6, 7, 8].
-Canonical approaches include diffusion- and patch-based algorithms, whose core principle involves leveraging neighborhood information from intact regions to fill missing areas via propagation or copy-paste mechanisms.
+Image inpainting is a fundamental computer vision task that restores missing or corrupted image regions.
+These imperfections can result from occlusions, sensor noise, data transmission loss, intentional editing, or physical damage.
+This capability makes it vital for applications like cultural preservation, photography, medical imaging, visual effects, and autonomous systems.
+However, achieving accurate and semantically coherent inpainting remains challenging.
+Conventional inpainting methods primarily rely on intrinsic statistical properties and local structural information of images [1, 2, 3, 4, 5, 6, 7, 8], whose core principle involves leveraging neighborhood information from intact regions to fill missing areas via propagation or copy-paste mechanisms.
 While adequate for small-scale, texturally regular regions, these methods exhibit significant limitations when confronted with extensive missing areas, complex structural boundaries, or semantically rich regions.
 They often generate artifacts such as blurred textures, fractured structural continuity, and semantically inconsistent content that undermines global image context.
-These limitations significantly constrain their effectiveness in real-world, high-fidelity imaging applications.
 The rapid development of deep learning—particularly Convolutional Neural Networks (CNNs), Generative Adversarial Networks (GANs), and emerging Transformer- and Diffusion-based models—has driven breakthroughs in complex pattern learning and high-fidelity image synthesis, thereby catalyzing a paradigm shift in image inpainting [9, 10, 11, 12, 13, 14, 15].
@@
 These models reconstruct intricate textural details while maintaining plausibility.
-For example, restored objects exhibit geometrically consistent structures, physically coherent illumination, and contextually appropriate depth relationships.
-State-of-the-art models can generate diverse yet equally plausible results, demonstrating performance that substantially surpasses conventional methods.
-Despite these advances and their widespread deployment, researchers still face persistent and critical challenges.
-A pivotal yet often overlooked issue is the inefficiency and inaccuracy stemming from informational redundancy.
-In many real-world applications, inpainting is fundamentally object-centric rather than indiscriminately processing the entire frame.
-Under such conditions, prevailing whole-image inpainting models — encompassing nearly all contemporary deep learning methods — incur substantial computational overhead by processing redundant background information.
-For instance, autonomous vehicles equipped with 8K cameras can generate up to 1.5 TB/s of data, with nearly 90% consisting of redundant background or irrelevant content; military reconnaissance satellites transmit massive raw imagery, with less than 0.1% containing actionable intelligence; and in medical X-ray imaging and industrial non-destructive testing, conventional area-array detectors achieve less than 0.13% efficiency in extracting useful information.
+Despite these advances, significant challenges remain.
+A key overlooked issue is inefficiency from information redundancy.
+In practice, inpainting is often object-centric rather than holistic.
+Consequently, most existing methods process redundant background, incurring substantial overhead.
+For example, across applications like autonomous driving, reconnaissance, and medical imaging, typically over 90% of data is background rather than target information.
 Whole-image models consume substantial computational resources analyzing redundant background information, leading to severe resource wastage and reduced accuracy in target-object reconstruction due to background interference, calling for a paradigm shift from holistic processing to targeted, object-focused inference.
-In confronting these challenges, biological vision systems offer an elegant and efficient solution.
-A core component of biological vision is the attention mechanism, which prioritizes task-relevant targets while suppressing redundant background information.
-This enables efficient, accurate, and energy-optimized information processing for target recognition and imaging in cluttered or dynamic scenes.
-This “object-first perception” strategy offers a promising solution for addressing low information density and background noise in inpainting: an inpainting pipeline could first localize and model target-object structures while excluding irrelevant background as interference.
-Such an approach would enhance imaging focus, computational efficiency, and output quality.
-Building on this analysis of the current state of inpainting, its technical bottlenecks, and biological inspiration, we propose a novel front-end paradigm: Image Inpainting via Visual-Neural-Inspired Specific Object-of-Interest Imaging technology(SIOI).
-Inspired by biological “object-first perception”, we propose a paradigm shift: Specific Object-of-Interest Imaging (SIOI).
-Unlike traditional methods, SIOI decouples the ‘what’ and ‘where’ of the object from the ‘how’ of its inpainting.
-It acts as a generic front-end that extracts a clean object representation, suppressing background noise, and can be plugged into any existing inpainting model.
-Our framework decomposes the imaging process into two distinct stages.
-Stage 1 focuses exclusively on extracting and modeling the structural information of target objects.
-This stage constitutes a dedicated information-extraction process designed to maximize target-relevant features while minimizing or eliminating background interference.
-Stage 2 performs coarse structural restoration guided by refined priors to support detail synthesis, refines the restoration, ensuring visual coherence and semantic consistency in the final output.
-The key innovation lies in Stage I; Stage II is deliberately kept simple in our experiments to prove the concept.
-By frontloading SIOI, the pipeline significantly reduces background distractions and concentrates structural modeling on critical target regions.
-This substantially reduces the data volume for processing, thereby improving computational efficiency while enhancing structural continuity and semantic accuracy.
+In confronting these challenges, biological vision offers an efficient solution.
+Visual attention mechanisms prioritize relevant targets while suppressing background, enabling efficient target recognition in complex scenes.
+This ”object-first” approach addresses information sparsity by first localizing targets while excluding background interference.
+This improves focus, efficiency, and output quality.
+Based on this analysis, we propose Specific Object-of-Interest Imaging (SIOI).
+SIOI decouples object identification from inpainting execution.
+It serves as a generic front-end that extracts clean object representations while suppressing background, pluggable into any inpainting model.
+Our framework employs two stages: (1) object perception, and (2) object inpainting.
+SIOI reduces background interference and focuses processing on target regions, improving efficiency and accuracy.
 As shown in Tab.
-1, unlike other image-restoration methods, this bio-inspired approach aligns with the hierarchical modeling of the visual cortex in neural networks, ensuring that computational efficiency parallels human attentional selection.
+1, this approach aligns with visual cortex hierarchy, mimicking human attentional mechanisms.
 Our principal contributions are:
   1. 1.
-    We introduce Specific Object-of-Interest Imaging (SIOI), a novel front-end paradigm for object-centric image processing.
-SIOI explicitly decouples target-object extraction from texture synthesis, providing a generic mechanism to reduce background interference and generate powerful structural priors for any subsequent task.
+    We propose Specific Object-of-Interest Imaging (SIOI), a novel framework that decouples object extraction from texture synthesis to reduce background interference and generate structural priors for subsequent tasks.
   2. 2.
-    We conduct a preliminary exploration of correlations between human vision and deep networks, integrating neuroscientific and psychological insights on hierarchical visual cortex processing to establish theoretical foundations.
+    We investigate correlations between human vision and deep networks, integrating neuroscientific insights from visual cortex hierarchy to establish theoretical foundations.
   3. 3.
-    We demonstrate that SIOI is an effective, plug-and-play front-end module.
-Through extensive experiments, we show that integrating SIOI consistently enhances the performance of various mainstream inpainting models across multiple metrics and datasets, while significantly improving robustness in extreme scenarios.
+    We demonstrate SIOI’s effectiveness as a plug-and-play module through comprehensive experiments, showing consistent performance improvements across metrics, datasets, and extreme scenarios.
 Table 1: Theoretical Framework Comparison
 Model Theoretical Framework Core Mechanism Processing Unit Biological Plausibility
-Traditional Digital image processing and classical machine learning techniques Handcrafted features and algorithms Entire image Not considered
-Deep-Learning Data-driven approaches enabled by deep neural networks Learned feature representations Entire image Not considered
-Ours Bio-inspired two-stage framework Object perception → Object inpainting Specific Object-of-Interest V1-V4 hierarchical simulation
+Traditional Classical image processing Handcrafted features Entire image Not considered
+Deep-Learning Deep neural networks Learned representations Entire image Not considered
+Ours Bio-inspired Object perception → inpainting Specific Object-of-Interest V1-V4 simulation
 2 Related Works
-Image inpainting is a long-studied and widely applied problem whose methods have undergone substantial paradigm shifts.
-From early mathematical models to modern deep generative architectures.
-While these approaches differ in implementation, they overwhelmingly share a whole-image processing paradigm.
+Image inpainting has evolved through substantial paradigm shifts, from early mathematical models to modern deep generative architectures.
+While these approaches differ in implementation, these methods predominantly employ whole-image processing.
 2.1 Traditional Mathematical Algorithms: Digital Image Processing and Machine Learning
-This period emphasized leveraging low-level statistical properties, local geometric structures, and hand-crafted mathematical models.
-Representative approaches include: 1) PDE- and diffusion-based models [1, 16, 17], which use isophote directions or engineered diffusion tensors to propagate color and gradient information from intact regions into missing areas while preserving edge coherence; and 2) patch-based methods [18, 19, 20], which exploit non-local self-similarity by searching for patches in intact regions that best match neighborhoods around missing regions and iteratively copying them to fill holes.
-Traditional mathematical approaches [21, 22, 23] reconstruct missing areas by propagating local structures or copying similar patches.
-They perform reasonably on small, regular textures but break down under large occlusions or semantically rich regions, since they lack object-level understanding.
+This period emphasized low-level statistics, local geometric structures, and hand-crafted models.
+Representative approaches include: 1) PDE- and diffusion-based models [1, 16, 17] that propagate color and gradient information while preserving edges; and 2) patch-based methods [18, 19, 20] that exploit non-local self-similarity by matching and copying patches from intact regions.
+These approaches [21, 22, 23] reconstruct missing areas through local propagation or patch copying.
+They perform well on small regular textures but struggle with large occlusions or semantically rich regions due to limited object-level understanding.
 2.2 Data-Driven Neural Network Approaches
-Deep-learning approaches substantially improve structural coherence and realism by learning high-level priors.
-Predominant paradigms include: (1) CNN encoder–decoder architectures [24, 4, 25, 26, 27], which exploit multi-level encoder features to progressively generate missing content in decoders, attention mechanisms and gated convolutions further improve focus on relevant regions; (2) GAN-based methods [28, 29, 30, 31, 32, 33, 34], which employ adversarial training between generators and discriminators to produce highly realistic outputs and excel at fine texture synthesis and alignment with natural image distributions; (3) Transformer-based approaches [4, 35, 36, 37, 38, 39, 40], which leverage self- and cross-attention to model long-range dependencies between patches and thereby capture global semantics and structural consistency, improving performance on complex scenes and large corruptions; and (4) diffusion models [41, 42, 43], which cast generation as a progressive denoising process that iteratively recovers image details from noise, offering fine-grained detail control, diverse outputs, and high fidelity.
+Deep learning methods improve structural coherence and realism by learning high-level priors.
+Key paradigms include: (1) CNN encoder-decoder architectures [24, 4, 25, 26, 27] using multi-level features to generate content, with attention mechanisms enhancing focus; (2) GAN-based approaches [28, 29, 30, 31, 32, 33, 34] using adversarial training to produce realistic outputs and fine textures; (3) Transformer-based methods [4, 35, 36, 37, 38, 39, 40] leveraging self- and cross-attention to model long-range dependencies and capture global semantics; and (4) Diffusion models [41, 42, 43] employing progressive denoising to recover image details with fine-grained control and high fidelity.
 Yet, most still operate on the entire image, wasting computation on irrelevant background and suffering from interference when target and background overlap.
@@
 Whole-image models thus suffer from several issues in object-centric tasks: 1) information redundancy and computational inefficiency — processing the entire image wastes computation on irrelevant background content, exacerbating the low effective information density typical in domains such as autonomous driving and satellite imaging; 2) background interference — complex or misleading backgrounds can corrupt the reconstruction of target object structure and semantics, especially when target and background share similar textures or colors; and 3) lack of target-specific guidance and focus — current methods rarely model explicitly which object should be targeted and therefore struggle to separate target signals from background noise during inpainting.
-Although some methods adopt user sketches or bounding boxes as guidance, these require extra interaction and still struggle to achieve precise target separation in complex scenes.
-Thus, a biologically inspired, target-focused imaging paradigm — one that actively suppresses irrelevant background information — is imperative for object-centric tasks.
-Our proposed method, Inpainting via Specific Object-of-Interest Imaging , directly addresses this gap.
-It introduces a front-loaded Specific Object-of-Interest Imaging module that algorithmically maximizes target-relevant information while suppressing background interference.
-This yields highly refined, target-centric signals for subsequent inpainting stages and thereby mitigates information redundancy and background interference.
+Although some methods adopt user sketches or bounding boxes as guidance[44, 45, 46], these require extra interaction and still struggle to achieve precise target separation in complex scenes.
+Therefore, a biologically inspired, target-focused imaging paradigm that actively suppresses irrelevant background information is essential for object-centric tasks.
+Our method, Inpainting via Specific Object-of-Interest Imaging, directly addresses this gap.
+It introduces a front-end Specific Object-of-Interest Imaging module that algorithmically maximizes target-relevant information and suppresses background interference.
+This produces refined target-centric signals for subsequent stages, mitigating information redundancy and background interference.
 2.3 Insights from Human Visual Mechanisms for Image imaging
-Psychologists and neuroscientists have reported direct correlations between deep neural networks (DNNs) and hierarchical processing in the visual cortex [44, 45].
-Human vision follows a hierarchical pathway: retina $\rightarrow$ lateral geniculate nucleus (LGN) $\rightarrow$ V1 (edge extraction) $\rightarrow$ V4 $\rightarrow$ IT (semantic processing) [46, 47].
-This observation motivates multi-stage imaging networks [48, 49], in which early convolutional layers focus on local structure repair (analogous to V1 edge detection), while deeper layers integrate object identity and scene context (analogous to IT semantic processing); skip connections can be viewed as a crude analogue of cortical feedback.
-Predictive-coding theory [50, 51] proposes that the brain leverages prior knowledge (e.g., Gestalt principles) to generate predictions and correct sensory errors.
-This idea parallels the generator–discriminator framework in deep inpainting [24, 4]: generators predict missing content by learning data distributions, while discriminators enforce semantic plausibility via adversarial training [52].
-At the level of attentional regulation, bottom-up attention captures salient features (e.g., high-frequency details) via thalamic structures such as the pulvinar [53, 54]; this motivates algorithmic mechanisms such as saliency-map-weighted losses that emphasize abrupt edge regions [55].
-Conversely, top-down attention involves prefrontal-cortex (PFC) modulation of visual responses according to task goals [54].
-This inspires task-oriented strategies in models: for example, conditional vectors can steer generation to enhance discriminative feature repair (e.g., facial keypoint reconstruction) and to improve global structural consistency for scene reconstruction.
-Examples of biologically inspired techniques include contour interpolation via edge-consistency losses [56], motivated by V1/V2 responses to illusory contours [57, 58, 59].
-Semantic guidance has been realized by mapping IT-like object activations.
-For instance, CLIP-guided diffusion has been used to align textual concepts with image generation [42].
-Memory-augmented imaging methods simulate hippocampal–cortical episodic recall by retrieving reference segments from external memory banks [60].
-Consequently, advancing image inpainting benefits from drawing principled inspiration from human visual mechanisms to identify promising directions for improvement.
-This ‘object-first’ perception allows efficient suppression of background noise and allocation of resources to relevant targets.
-Such mechanisms directly motivate our proposed Specific Object-of-Interest Imaging, which introduces an analogous front-end object-first module into the inpainting pipeline.”
-In summary, while existing methods have advanced in realism and fidelity, they all share a whole-image bias that is ill-suited for object-centric tasks.
-Inspired by biological vision, we propose a new front-end paradigm—Specific Object-of-Interest Imaging—to fill this critical gap.
+[…]
+In summary, despite advances in realism, existing methods retain a whole-image bias ill-suited for object-centric tasks.
+Inspired by biological vision, we propose Specific Object-of-Interest Imaging to address this gap.
 3 Proposed Methodology
-This section delineates our proposed framework for Image Inpainting.
-Our approach introduces a novel paradigm for this task.
-On a conceptual level, it shares a similar goal—focusing on a specific object region—with other imaging techniques, such as [61, 62, 63, 64, 65], but the underlying methodology is uniquely our own work.
+This section presents our novel image inpainting framework.
+While conceptually sharing the object-focus goal with other techniques [63, 64, 65, 66, 67], our methodology is fundamentally distinct.
 The processing pipeline is illustrated in Fig.
@@
 3.1 Specific Object-of-Interest Imaging
-Specific Object-of-Interest Imaging is also known as Attention-Targeted Imaging.
-Desimone and Duncan’s competitive-bias model [66] posits that multiple objects compete for activation resources in the visual cortex.
-Attention serves as a bias signal that preferentially amplifies target objects and suppresses background information.
-This provides a neuroscientific foundation for implementing background suppression and object-enhancement saliency selection in computational networks.
+Also termed Attention-Targeted Imaging, SIOI builds on neuroscientific foundations.
+Desimone and Duncan’s model [68] posits visual cortex competition among objects for activation resources.
+Attention acts as a bias signal amplifying targets while suppressing background, providing a neuroscientific basis for computational background suppression and object enhancement.
 O’Craven et al.
-[67] demonstrated with fMRI that objects act as fundamental attentional units, exhibiting holistic enhancement even when component features are spatially segregated.
-Other studies [68, 69] further showed that object boundaries automatically attract attention, recruiting V1–V4 regions and intensifying internal feature encoding.
+[69] fMRI studies show objects as fundamental attentional units with holistic enhancement.
+Studies [70, 71] show object boundaries automatically attract attention, engaging V1-V4 regions.
 Perry et al.
-[70] reported that foveal cortical magnification increases encoding density, and that task-dependent plasticity enables high-resolution perception of behaviorally relevant regions.
-Carrasco and Barbot [71, 72] systematically characterized attention’s dual regulation of spatial resolution: exogenous attention enhances frequency sensitivity at fixation loci, whereas endogenous attention reallocates resolution according to task demands.
-Barbot and Carrasco further showed that attentional guidance amplifies high-frequency recognition, providing a basis for region-adaptive reconstruction.
+[72] reported foveal magnification increases encoding density and task-dependent plasticity enables high-resolution perception.
+Carrasco and Barbot [73, 74] characterized attention’s dual regulation: exogenous enhances frequency sensitivity, endogenous reallocates resolution.
+They further showed attentional guidance amplifies high-frequency recognition, enabling region-adaptive reconstruction.
 Ekman et al.
-[73] demonstrated that attentional signals propagate along object boundaries in V1, inducing internal feature enhancement — a phenomenon translatable to frequency-domain focusing mechanisms in networks (e.g., concentrating computational resources using object masks).Collectively, these findings reveal an object-based prioritization mechanism: the visual system preferentially allocates resources to enhance detailed representation of target objects.
-These observations directly inspired our object-detection–guided saliency-enhancement branch, which applies high-frequency component weighting to target regions.
-We propose an approach that integrates saliency modeling with biologically inspired, object-guided image imaging to construct a cognitively plausible, interpretable, and task-relevant image inpainting system (see Fig.
+[75] showed attentional signals propagate along V1 object boundaries, inducing feature enhancement—translating to network frequency-domain focusing.
+Collectively, these reveal object-based prioritization: preferential resource allocation to enhance target representation.
+These directly inspired our object-detection-guided saliency enhancement applying high-frequency weighting to targets.
+We integrate saliency modeling with bio-inspired object guidance to construct a cognitively plausible inpainting system (Fig.
 2).
 Figure 2: Image inpainting framework based on the visual neural Specific Object-of-Interest Imaging.
-Neuroanatomically, visual processing follows a hierarchical architecture: local feature detection (retina) $\rightarrow$ spatial integration (LGN) $\rightarrow$ orientation selectivity (V1) $\rightarrow$ holistic object representation (IT cortex).
-This multi-stage processing implies that a natural-scene image $k$ can be viewed as a superposition of multiple latent target objects, with attentional modulation in V4 amplifying target-specific features while suppressing background information.
-Grounded in neuronal population-coding theory and Hebbian plasticity, we formalize this composition as a linear superposition model:
+Anatomically, the visual system processes information hierarchically: local feature detection (retina) $\rightarrow$ spatial integration (LGN) $\rightarrow$ orientation selectivity (V1) $\rightarrow$ holistic object representation (IT cortex).
+This hierarchical processing suggests that a natural image $k$ comprises multiple latent objects, where V4-mediated attentional modulation enhances target features and suppresses background information.
+Based on neuronal population-coding and Hebbian plasticity, we model this composition through linear superposition:
 $k=\sum_{i=1}^{M}{\{}k_{i}\}$ (1)
-where $k\in\mathbb{R}^{H\times W\times C}$ denote the scene tensor, $k_{i}\in\mathbb{R}^{h_{i}\times w_{i}\times C}$ the $i$-th target object, and $M\in\mathbb{Z}^{+}$ the total number of objects.
-Each $k_{i}$ corresponds to a semantically complete entity in the scene (e.g., pedestrians or vehicles).
-Amodal content completion emphasizes restoring full object structures in occluded regions.
-The core objective is anatomically precise object separation: accurately extracting specific targets $k_{i}$ ($i\in\{1,\dots,M\}$) from complex backgrounds.
-This parallels the “cocktail-party” effect in human perception when attending to a target in cluttered scenes.
-To overcome the representational fragility of traditional methods in noisy environments, we propose a biologically inspired dynamic feature-modulation paradigm.
-Its mathematical core is a target-sensitive measurement matrix $\Gamma$.
-Via a differentiable learning mechanism, $\Gamma$ induces a cross-scale feature set $\{\xi_{j}\}_{j=1}^{N}$, reflecting the hierarchical abstraction from V1 simple cells (edge/texture responses) to IT complex cells (holistic shape responses).
-Below we make these notions precise.
-For a target index $\mu\in\{1,\dots,M\}$, we write
+where $k\in\mathbb{R}^{H\times W\times C}$ is the scene tensor, $k_{i}\in\mathbb{R}^{h_{i}\times w_{i}\times C}$ represents the $i$-th target object, and $M\in\mathbb{Z}^{+}$ is the total object count.
+Each $k_{i}$ represents a semantically coherent entity (e.g., a pedestrian or vehicle).
+Amodal completion focuses on reconstructing fully occluded object structures.
+The core objective is precise object separation—extracting specific targets $k_{i}$ ($i\in\{1,\dots,M\}$) from cluttered backgrounds.
+This mirrors the “cocktail-party” effect, where humans focus on a target speaker amidst noise.
+To overcome the limitations of traditional methods in noise, we introduce a bio-inspired dynamic feature modulation approach.
+Its mathematical basis is a target-sensitive measurement matrix $\Gamma$.
+Through differentiable learning, $\Gamma$ produces a multi-scale feature set $\{\xi_{j}\}_{j=1}^{N}$, capturing the hierarchy from V1 simple cells (edge/texture) to IT complex cells (shape representation).
+We formalize these concepts next.
+For a target index $\mu\in\{1,\dots,M\}$, we have:
 $\Gamma k_{\mu}=\sum_{j=1}^{N}\omega_{j,\mu}\xi_{j},$ (2)
-where $\Gamma\in\mathbb{R}^{N\times D}$ denotes the measurement matrix, $\xi_{j}\in\mathbb{R}^{D}$ the feature-basis vectors, and $\omega_{j,\mu}\in\mathbb{R}$ the feature weights associated with target $\mu$.
-When optimization drives the object- and target-specific weights $\omega_{j,\mu}$ to concentrate on the feature patterns of a particular $\mu$, the system exhibits a selective-gain effect: scene measurements become dominated by target-related information flow.
-These dynamics are expressed as
-$\Gamma k\;=\;\begin{bmatrix}\sum_{m=1}^{M}\eta_{1,m}\,\xi_{1}\\[4.0pt] \vdots\\[4.0pt] \sum_{m=1}^{M}\eta_{N,m}\,\xi_{N}\end{bmatrix}\approx\widehat{k}_{\mu},$ (3)
-where $\eta_{j,m}$ denotes the contribution of object $m$ to feature channel $j$.
-Measurements concentrate on the specific target when $\eta_{j,m}\to\delta_{m\mu}$ for each feature channel $j$, i.e., when the contributions from non-target objects vanish.
-Functionally, this formulation mirrors the hierarchical transformation from V1 simple cells (edge/texture responses) to IT complex cells (holistic shape responses), with $\Gamma$ playing the role of attentional gain control in V4.
-V1 encodes local orientations, while V4 integrates global contours under top-down modulation, consistent with our FFC-based frequency-domain focusing.
-From a functional-analysis perspective, this process can be interpreted as a target-guided modulation on an appropriate Riemannian manifold:
+where $\Gamma\in\mathbb{R}^{N\times D}$ is the measurement matrix, $\xi_{j}\in\mathbb{R}^{D}$ are the basis vectors, and $\omega_{j,\mu}\in\mathbb{R}$ are the target-specific weights.
+When optimized, the weights $\omega_{j,\mu}$ concentrate on patterns for $\mu$, creating a selective-gain effect where scene measurements are dominated by the target.
+This leads to the scene-level measurement:
+$\Gamma k=\begin{bmatrix}\sum_{m=1}^{M}\eta_{1,m}\xi_{1}\\ \vdots\\ \sum_{m=1}^{M}\eta_{N,m}\xi_{N}\end{bmatrix}\approx\widehat{k}_{\mu},$ (3)
+where $\eta_{j,m}$ is the contribution of object $m$ to channel $j$. Measurements focus on the target when $\eta_{j,m}\to\delta_{m\mu}$ for all $j$, meaning non-target contributions vanish.
+Functionally, this models the transformation from V1 (edges/textures) to IT (shapes), with $\Gamma$ acting as an attentional gain in V4.
+V1 encodes local orientations, while V4 integrates global contours, aligning with our frequency-domain (FFC) focus.
+From a functional view, this process is a target-guided modulation on a Riemannian manifold:
 $\Gamma k=\sum_{m=1}^{M}\Gamma\omega_{m}\otimes k_{m}\approx\hat{k}^{\prime}_{0}$ (4)
-where $\otimes$ denotes the diffeomorphism operator in the feature space.
-Establishing a rigorous mathematical model is essential for precise object separation.
-Consider an input scene containing $N$ latent targets.
-Sparse representation of individual objects.
-Each object $k_{i}$ ($i=0,1,\dots,N-1$) can be represented sparsely in a high-dimensional dictionary basis $\Psi$:
+where $\otimes$ is the diffeomorphism operator in feature space.
+A rigorous mathematical model is essential for precise object separation.
+For a scene with $N$ latent targets, each object admits a sparse representation.
+Each object $k_{i}$ ($i=0,1,\dots,N-1$) is sparsely represented in a high-dimensional basis $\Psi$:
 $k_{i}=\sum_{j=1}^{M}\psi_{ij}x_{ij}$ (5)
-where $\psi_{ij}$ denotes the adaptive sparse basis and $x_{ij}$ the sparse coefficients.
-Panoramic scene decomposition indicates that the holistic scene can be expressed as a linear combination of objects:
+where $\psi_{ij}$ is the adaptive basis and $x_{ij}$ the sparse coefficients. The scene decomposes into a linear combination of objects:
 $k=\sum_{i=0}^{N-1}k_{i}=\sum_{i=0}^{N-1}\sum_{j=1}^{M}\psi_{ij}x_{ij}$ (6)
-subject to the following constraints: $\|\mathbf{x}_{i}\|_{0}\ll M\quad\text{(Sparsity)},\quad\psi_{ij}\in\mathbb{R}^{D}\quad\text{(Dictionary Basis)},\quad x_{ij}\in\mathbb{R}\quad\text{(Sparse Coefficient)}.$ This representation satisfies the principles of compressed sensing, thereby ensuring object separability in the feature space.
-This theoretical model is implemented through a bio-inspired neural architecture.
-The Feedforward Feature Selection Network simulates the ventral (“what”) visual pathway by performing target-sensitive filtering on masked defective inputs $M_{h}$.
-The Target Reconstruction Network mimics the spatial integration function of the parietal cortex and is dedicated to target reconstruction.
-The training process is formulated as a physically constrained optimization problem:
+subject to sparsity ($\|\mathbf{x}_{i}\|_{0}\ll M$), basis constraints ($\psi_{ij}\in\mathbb{R}^{D}$), and coefficient constraints ($x_{ij}\in\mathbb{R}$).
+This compressed sensing formulation ensures object separability in feature space.
+We implement this via a bio-inspired neural architecture.
+The Feedforward Feature Selection Network mimics the ventral (”what”) pathway, filtering masked inputs $M_{h}$.
+The Target Reconstruction Network emulates parietal spatial integration for target reconstruction.
+Training solves a constrained optimization problem:
 $\{\xi,\theta\}=\arg\min\frac{1}{H}\sum_{h=1}^{H}\mathcal{C}(N_{h},\tilde{N}_{h})$ (7)
-where $\mathcal{C}(\cdot)$ denotes the loss function. Upon convergence, the FECNet weights $\xi$ encode target knowledge into a feature filter $\phi$. Applying $\phi$ to the panoramic scene yields:
+where $\mathcal{C}(\cdot)$ is the loss function. At convergence, FECNet weights $\xi$ form a feature filter $\phi$. Applying $\phi$ gives:
 $\phi k=\phi\left(\sum_{i=0}^{N-1}\sum_{j=1}^{M}\psi_{ij}x_{ij}\right)=\sum_{j=1}^{M}\psi_{0j}x_{0j}=\hat{k}_{0}$ (8)
-Since $\phi$ is a linear operator, it satisfies the following projection property:
+As a linear operator, $\phi$ has the projection property:
 $\phi(x_{ij})=\begin{cases}x_{0j},&\text{if }i=0\\ 0,&\text{if }i\neq 0\end{cases}$ (9)
-Assume the dictionary $\Psi$ satisfies the Restricted Isometry Property (RIP), i.e.
-$\mathrm{RIP}(\Psi)\leq\delta_{2s}$ (separability condition).
-The underlying mechanism is that the filter $\phi$ implements an (approximate) orthogonal complement in feature space, $\phi=I-P_{\mathrm{bg}}$, thereby suppressing non-target components (i.e., those with $i\neq 0$).
-A closed-form estimator for $\phi$ can be obtained by minimizing a least-squares criterion and expressed via the Moore–Penrose pseudoinverse.
-Defining $k^{(h)}$ as the vectorized scene for sample $h$ and $\widehat{k}_{0}^{(h)}$ as its corresponding target component, we may write
+Assuming the dictionary $\Psi$ satisfies the Restricted Isometry Property (RIP), i.e., $\mathrm{RIP}(\Psi)\leq\delta_{2s}$ (separability condition), the filter $\phi$ approximates an orthogonal complement in feature space ($\phi=I-P_{\mathrm{bg}}$), suppressing non-target components ($i\neq 0$).
+A closed-form estimator for $\phi$ is derived by minimizing a least-squares criterion using the Moore-Penrose pseudoinverse:
 $\phi=\arg\min_{\phi^{\prime}}\sum_{h}\Big\|\hat{k}_{0}^{(h)}-\phi^{\prime}\Big(\sum_{i=0}^{N-1}k_{i}^{(h)}\Big)\Big\|_{F}^{2}$ (10)
 $\Rightarrow\quad\phi=\Big(\sum_{h}\hat{k}_{0}^{(h)}\,k^{(h)\,T}\Big)\Big(\sum_{h}k^{(h)}k^{(h)\,T}\Big)^{\dagger}$
-where denotes the Moore–Penrose pseudoinverse, and $k^{(h)}\in\mathbb{R}^{D\times 1}$ is the vectorized scene representation for sample $h$.
-We further constrain $\phi=\mathbf{I}-\mathbf{P}_{\mathrm{bg}}$ as an (approximate) orthogonal projection onto the target subspace.
-Equivalently, $\phi$ may be viewed as the best approximation to the ideal projection $I-P_{\mathrm{bg}}$ in a suitable Hilbert space $\mathcal{H}$:
+where denotes the Moore-Penrose pseudoinverse, and $k^{(h)}\in\mathbb{R}^{D\times 1}$ is the vectorized scene for sample $h$.
+We constrain $\phi=\mathbf{I}-\mathbf{P}_{\mathrm{bg}}$ as an approximate orthogonal projection onto the target subspace.
+Equivalently, $\phi$ is the best approximation to $I-P_{\mathrm{bg}}$ in a Hilbert space $\mathcal{H}$:
 $\phi^{\star}=\arg\min_{\phi\in\mathcal{H}}\;\big\|\,\phi-(I-P_{\mathrm{bg}})\,\big\|_{F}.$ (11)
-Here $\mathcal{H}$ is a Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and induced norm $\|\cdot\|$; $\|\cdot\|_{F}$ denotes the Frobenius norm for matrices, $\|A\|_{F}=\sqrt{\sum_{i}\sum_{j}|a_{ij}|^{2}}$.
-Let $\mathbf{q}_{xy}$ and $\mathbf{k}_{ij}$ denote learned query and key feature vectors at spatial locations $(x,y)$ and $(i,j)$, respectively, and let $d$ be the feature dimensionality used for softmax scaling.
-Under bio-inspired attentional modulation we compute attention weights by
+Here $\mathcal{H}$ is a Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and norm $\|\cdot\|$; $\|\cdot\|_{F}$ is the Frobenius norm, where $\|A\|_{F}=\sqrt{\sum_{i}\sum_{j}|a_{ij}|^{2}}$.
+Let $\mathbf{q}_{xy}$ and $\mathbf{k}_{ij}$ be learned query and key vectors at locations $(x,y)$ and $(i,j)$, respectively, with $d$ as the feature dimensionality for softmax scaling.
+Under bio-inspired attentional modulation, attention weights are computed as:
 $\alpha_{xy}=\frac{\exp\!\big(\langle\mathbf{q}_{xy},\mathbf{k}_{ij}\rangle/\sqrt{d}\big)}{\sum_{i,j}\exp\!\big(\langle\mathbf{q}_{xy},\mathbf{k}_{ij}\rangle/\sqrt{d}\big)},$ (12)
-where the denominator normalizes over all spatial locations $(i,j)$.
-When $\mathbf{q}_{xy}$ and $\mathbf{k}_{ij}$ are highly similar, the corresponding attention weight $\alpha_{xy}$ approaches one, consistent with Feature Integration Theory: biased competition in V4 enhances target features via attentional gain.
-Require that the Kullback–Leibler divergence between the target and background distributions satisfies
+where the denominator normalizes over all locations $(i,j)$.
+When $\mathbf{q}_{xy}$ and $\mathbf{k}_{ij}$ are similar, $\alpha_{xy}$ approaches 1, consistent with Feature Integration Theory: biased competition in V4 enhances targets via attentional gain.
+We require the Kullback-Leibler (KL) divergence between target and background distributions to satisfy:
 $D_{\mathrm{KL}}\big(P_{\mathrm{target}}\|P_{\mathrm{background}}\big)>\sum_{i=2}^{M}D_{\mathrm{KL}}\big(P_{\mathrm{target}}\|P_{\mathrm{object}_{i}}\big),$ (13)
-where $D_{\mathrm{KL}}(P\|Q)=\sum_{i}P_{i}\log\frac{P_{i}}{Q_{i}}$ denotes the Kullback–Leibler divergence.
-By large-deviation principles, when the distributional discrepancy between the target and background dominates inter-object interference, empirical distributions concentrate on distributions minimizing relative entropy; consequently, the estimator tends to concentrate on the target distribution, which implies reduction of the estimated target entropy under suitable conditions.
-Hence, under these assumptions there exists a unique solution $\widehat{k}_{0}$ such that
+where $D_{\mathrm{KL}}(P\|Q)$ is the KL divergence.
+By large-deviation principles, if the target-background discrepancy dominates inter-object interference, empirical distributions concentrate on entropy-minimizing distributions.
+Thus, the estimator concentrates on the target distribution, reducing estimated target entropy under suitable conditions.
+Hence, under these assumptions, a unique solution $\widehat{k}_{0}$ exists such that:
 $\lim_{t\to\infty}H\big(\widehat{k}_{0}^{(t)}\big)=H_{\mathrm{target}},$ (14)
-where $H(X)=-\sum_{x\in\mathcal{X}}p(x)\log p(x)$ denotes the Shannon entropy.
-Under Assumptions A1–A3 (sparsity of object representations, the RIP on dictionary $\Psi$, and bounded measurement noise), there exists a linear operator $\phi$ (learnable by FECNet) that approximates the ideal target projector $I-P_{\mathrm{bg}}$.
-By compressed-sensing theory, if $\mathrm{RIP}(\Psi)\leq\delta_{2s}$ then the learned projector $\phi$ (approximating $I-P_{\mathrm{bg}}$) admits the error bound
+where $H(X)$ is the Shannon entropy.
+Under Assumptions A1–A3 (object sparsity, RIP on $\Psi$, and bounded noise), a linear operator $\phi$ (learnable by FECNet) approximates the ideal projector $I-P_{\mathrm{bg}}$.
+By compressed-sensing theory, if $\mathrm{RIP}(\Psi)\leq\delta_{2s}$, the learned projector $\phi$ admits the error bound:
 $\|\phi k-\widehat{k}_{0}\|_{2}\leq C_{1}\,\sigma_{s}(k)+C_{2}\,\epsilon,$ (15)
-where $C_{1}$ and $C_{2}$ depend on the RIP constants and the dictionary condition number; $\sigma_{s}(k)$ denotes the error of the best $s$-term approximation of $k$; and $\epsilon$ denotes an upper bound on the measurement noise.
-Together, these results provide a rigorous foundation for the proposed object-centric imaging approach.
-By casting target separation as a formal projection problem under sparsity constraints and employing information-theoretic concentration arguments to characterize convergence, we show that under the stated assumptions the Specific Object-of-Interest Imaging system provably isolates the target and converges to an optimal reconstruction that up to the approximation and noise terms given above.
+where $C_{1}$ and $C_{2}$ depend on RIP constants and the dictionary condition number; $\sigma_{s}(k)$ is the best $s$-term approximation error; and $\epsilon$ bounds the measurement noise.
+These results establish a rigorous foundation for our object-centric imaging approach.
+By formulating target separation as a sparsity-constrained projection problem and using information-theoretic arguments, we show that under our assumptions, the SIOI system provably isolates the target and converges to an optimal reconstruction, up to approximation and noise errors.
 3.2 Image Inpainting System
 Figure 3: Processing flow of the proposed image-inpainting system.
-To enhance structural coherence and reconstruction accuracy for occluded images, we propose a two-stage framework grounded in specific-object-of-interest imaging principles (Fig.
+We propose a two-stage framework based on specific-object-of-interest imaging to enhance structural coherence and reconstruction accuracy for occluded images (Fig.
 3).
-Departing from traditional pixel-level completion in RGB space, our architecture adopts an object-first perception inspired by biological vision, decomposing the task into structure-prioritized progressive refinements.
-Emphasizing semantically coherent reconstruction, we decompose the imaging pipeline into two stages: (1) object perception $\rightarrow$ (2) object inpainting.
-The core innovation is SIOI, which compels the model to focus on the target object while suppressing irrelevant background interference.
-This reduces computational overhead by concentrating structural modeling on critical regions (see Fig.
+Unlike traditional RGB-space pixel completion, our architecture adopts a biologically inspired object-first approach, decomposing the task into structure-prioritized refinements.
+We decompose the pipeline into: (1) object perception $\rightarrow$ (2) object inpainting, emphasizing semantic coherence.
+The core innovation, SIOI, directs model focus to the target while suppressing background interference.
+This reduces computation by concentrating on critical target regions (see Fig.
 4).
 Figure 4: Architecture of the proposed Specific Object-of-Interest Imaging.
-Stage I: Specific Object-of-Interest Imaging Biological vision begins with retinal photoreceptors converting incoming light into neural signals.
-Subsequent processing in the lateral geniculate nucleus (LGN) integrates spatial information and suppresses high-frequency noise via center–surround antagonism.
-Feature extraction proceeds in primary visual cortex (V1), where simple cells respond to oriented edges and spatial-frequency textures [46].
-Object prioritization amplifies target representations through bottom-up attentional biasing [66], suppressing background interference—functionally analogous to selective attention in the ”cocktail-party” effect.
-Our SIOI module computationally embodies this hierarchy.
-We process corrupted inputs with modified U-Net encoders to extract multi-scale features, mimicking V1 orientation selectivity.
-Embedding Fast Fourier Convolution (FFC) modules in multiple encoding and decoding processes captures long-range dependencies through spectral transformation, simulating attentional modulation in V4 that amplifies target signals.
-The spectral processing can be written as
+Stage I: Specific Object-of-Interest Imaging Biological vision starts with retinal phototransduction.
+Subsequent LGN processing integrates spatial information and suppresses high-frequency noise through center-surround antagonism.
+In primary visual cortex (V1), simple cells extract oriented edges and spatial-frequency textures [49].
+Object prioritization amplifies targets via bottom-up attention [68], suppressing background interference—analogous to the ”cocktail-party” effect.
+Our SIOI module computationally implements this hierarchy.
+Using modified Encoder-Decoder, we process corrupted inputs to extract multi-scale features, mimicking V1 orientation selectivity.
+FFC modules in encoding/decoding capture long-range dependencies via spectral transformation, simulating V4 attentional modulation that amplifies target signals.
+The spectral processing is:
 $\mathcal{F}_{\mathrm{freq}}=\mathrm{Conv}_{\mathrm{freq}}\big(\mathrm{FFT2}(\mathcal{F})\big),\qquad\mathcal{F}^{\prime}=\mathrm{IFFT2}(\mathcal{F}_{\mathrm{freq}}),$ (16)
-where $\mathrm{FFT2}$ and $\mathrm{IFFT2}$ denote the 2-D Fourier transform and its inverse, respectively, and $\mathrm{Conv}_{\mathrm{freq}}$ is a learnable convolutional operator applied in the frequency domain.(Fig.
-5)
-The output of the SIOI module is a preliminary, yet structurally coherent, object map contain local defects.
-As shown in Fig.
+where $\mathrm{FFT2}$$\mathrm{IFFT2}$ are 2-D Fourier transform/inverse, and $\mathrm{Conv}_{\mathrm{freq}}$ is a learnable frequency-domain convolutional operator (Fig. 4).
+The SIOI module outputs a preliminary but structurally coherent object map with local defects, as shown in Fig.
 4.
-Biologically, this parallels V1 edge detection and V4-mediated object-based attention, in which attentional gain enhances target representations while suppressing background activity [74].
-Our FFC modules enhance target-relevant spectral components while attenuating background spectra.
-Functionally, FFC operations correspond to V1-like orientation detection, whereas the learned attention weights emulate pulvinar–thalamic modulation of cortical processing.
-Consequently, Stage I produces structurally coherent object maps despite occlusions, providing essential priors for subsequent texture synthesis and refinement.This spectral processing overcomes the limitations of standard convolutions for modeling long-range dependencies and significantly improves structural consistency.
-Together, these mechanisms reduce information redundancy.
-The SIOI module serves as a generic, front-end ”visual processor” that transforms a complex, occluded scene into a purified, object-centric representation.
-It provides a powerful intermediate representation that is the cornerstone of our framework, explicitly containing the structural and semantic priors necessary for high-fidelity inpainting.
-Stage II: Object Inpainting (Coarse inpainting + Fine inpainting) To validate the efficacy of the SIOI module independently of advanced inpainting architectures, we deliberately employ a straightforward inpainting network in Stage II.
-Unlike conventional image inpainting methods, our model eliminates the need for additional manually annotated corruption masks and is capable of automatically identifying and inpainting the corrupted regions of the target.
-It is critical to emphasize that the innovation of this work is not this network itself, but rather the SIOI module.
-The purpose of Stage II is solely to demonstrate that the priors generated by SIOI are so potent that they enable even a simple network to achieve high-quality reconstruction.
-This stage is loosely analogous to the role of higher visual areas in integrating semantic information for detailed perception.
-Our Structural Recovery Module (SRM) takes the SIOI output and performs contour interpolation and shape completion using multi-scale dilated convolutions and feature-pyramid fusion:
+Biologically, this mimics V1 edge detection and V4-mediated attention, where attentional gain enhances targets and suppresses background activity [76].
+Our FFC modules amplify target-relevant spectral components and attenuate background spectra.
+Functionally, FFC operations resemble V1 orientation detection, while attention weights emulate pulvinar-thalamic cortical modulation.
+Thus, Stage I generates structurally coherent object maps despite occlusions, providing essential priors for texture synthesis and refinement.
+This spectral processing improves upon standard convolutions by modeling long-range dependencies and enhancing structural consistency.
+These mechanisms collectively reduce information redundancy.
+The SIOI module acts as a front-end visual processor, transforming complex, occluded scenes into purified, object-centric representations.
+It provides a key intermediate representation, containing structural and semantic priors essential for high-fidelity inpainting.
+Stage II: Object Inpainting (Coarse + Fine) To validate SIOI module efficacy independently, we use a simple inpainting network in Stage II.
+Unlike conventional methods, our model requires no manual masks and automatically identifies and inpaints target corrupted regions.
+Critically, the innovation is the SIOI module, not this network.
+Stage II demonstrates that SIOI priors enable high-quality reconstruction even with a simple network.
+This stage analogizes higher visual areas integrating semantic information for detailed perception.
+Our Structural Recovery Module (SRM) processes SIOI output with multi-scale dilated convolutions and feature-pyramid fusion for contour interpolation and shape completion:
 $\mathcal{F}_{\mathrm{MSDC}}=\mathrm{Concat}\left(\bigcup_{r\in\{1,2,4,6\}}\mathrm{Conv}_{r}(x)\right)$ (17)
-An FPN-like fusion integrates encoder–decoder hierarchical features, and subpixel-convolution upsampling is used to suppress checkerboard artifacts.
+FPN-like fusion integrates hierarchical features, and subpixel-convolution upsampling suppresses checkerboard artifacts.
 The output is a complete structural feature tensor $\mathcal{F}_{\mathrm{SRM}}\in\mathbb{R}^{64\times H\times W}$.
 Figure 5: Architecture of the Simple Object Inpainting.
-Functionally, both biological and computational systems achieve structural coherence: biological vision generates shape priors via predictive-coding [50], while the SRM fuses multi-scale context through MSDC and feature-pyramid fusion.
-This design helps to avoid texture blurring and semantic mismatches.
-Technically, MSDC operations correspond to V2–V3 contour integration, while feature-pyramid fusion approximates cortical feedback modulation.
-SRM enhances structural integrity in partially occluded object maps from Stage I, enabling a coarse-to-fine imaging strategy.
+Functionally, both systems achieve structural coherence: biological vision uses predictive-coding for shape priors [53], while SRM fuses multi-scale context via MSDC and feature-pyramid fusion.
+This design avoids texture blurring and semantic mismatches.
+Technically, MSDC corresponds to V2–V3 contour integration, and feature-pyramid fusion approximates cortical feedback.
+SRM enhances structural integrity in occluded maps from Stage I, enabling coarse-to-fine imaging.
 $\mathcal{F}_{\mathrm{SRM}}=\mathcal{R}(I_{\mathrm{coarse}}),\qquad\mathcal{F}_{\mathrm{SRM}}\in\mathbb{R}^{64\times H\times W},$ (18)
-where $\mathcal{R}(\cdot)$ denotes SRM processing and $I_{\mathrm{coarse}}$ denotes the coarsely restored map.
-Subsequently, the Global Detail Recovery Module (GDRM) utilizes a symmetric encoder-decoder with skip connections to synthesize fine-grained textures onto the recovered structure, guided by the robust priors from Stage I.
-Our computational counterpart, the Global Detail Recovery Module, implements a symmetric encoder–decoder architecture that consumes the 64-channel structural features $\mathcal{F}_{\mathrm{SRM}}\in\mathbb{R}^{64\times H\times W}$ produced by the SRM.
-This simple design choice underscores our core claim: The SIOI module is a versatile, plug-and-play component that can elevate the performance of any existing inpainting pipeline by providing superior object-specific guidance.
+where $\mathcal{R}(\cdot)$ is SRM processing and $I_{\mathrm{coarse}}$ is the coarsely restored map.
+The Global Detail Recovery Module (GDRM) then uses a symmetric encoder-decoder with skip connections to synthesize fine textures, guided by Stage I priors.
+GDRM implements a symmetric encoder-decoder that consumes the 64-channel features $\mathcal{F}_{\mathrm{SRM}}$ from SRM.
+This design underscores that SIOI is a versatile, plug-and-play component that elevates any inpainting pipeline via superior object-specific guidance.
 3.3 Loss function
-To achieve high-fidelity reconstruction of defective object images, we design a multi-branch fused loss that jointly promotes structural coherence, fine-detail reconstruction, and global visual-style consistency.
-The composite loss comprises three components: a weighted reconstruction loss [48, 75], a perceptual loss [76], and a style loss [77].
-Integrated with principled weighting and multi-stage optimization strategies, the loss ensures that synthesized images exhibit improved semantic structure and textural fidelity.
+To achieve high-fidelity reconstruction of defective object images, we propose a multi-branch composite loss function that simultaneously enhances structural coherence, fine-detail recovery, and global style consistency.
+It integrates three components: (1) a weighted reconstruction loss [51, 77], (2) a perceptual loss [78], and (3) a style loss [79].
+With carefully designed weighting schemes and a multi-stage optimization strategy, this composite loss ensures semantically coherent structures and texturally faithful details in the synthesized images.
 Weighted Reconstruction Loss.
-We introduce a spatially weighted reconstruction loss based on target maps.
+We first introduce a spatially weighted reconstruction loss guided by target-specific maps.
 An explicit weight map $w(x,y)$ differentially weights per-pixel reconstruction errors in the $L_{1}$$L_{2}$ terms:
@@
 where $\hat{I}(x,y)$ denotes the predicted intensity at pixel $(x,y)$, $I_{\mathrm{gt}}(x,y)$ the corresponding ground truth, $I_{i}^{\mathrm{pred}}$ and $I_{i}^{\mathrm{gt}}$ the RGB vectors at pixel $i$, $w_{i}$ the pixel-level weight, and $\alpha$ a scalar weighting factor (empirically set to 4).
-Higher weights are assigned to semantic target regions and lower weights to background regions; this directs the network to prioritize detail reconstruction in structurally critical areas and effectively mitigates edge blurring and structural discontinuities.
-This mechanism intensifies supervision in target regions while reducing gradient interference from irrelevant areas.
-Perceptual Loss. To compensate for pixel-level losses’ inability to capture high-level semantics, we incorporate a perceptual loss computed from feature activations of a pretrained VGG-16 network:
+This weighting scheme prioritizes detail reconstruction in semantically important regions while reducing emphasis on background areas, effectively mitigating edge blurring and structural discontinuities.
+This approach enhances supervision on target regions while minimizing interference from irrelevant background gradients.
+Perceptual Loss. To address the limitation of pixel-level losses in capturing high-level semantics, we incorporate a perceptual loss derived from feature activations of a pretrained VGG-16 network:
 $\mathcal{L}_{\mathrm{perc}}\;=\;\sum_{l\in\mathcal{L}}\frac{1}{C_{l}H_{l}W_{l}}\,\big\|\phi_{l}(I^{\mathrm{pred}})-\phi_{l}(I^{\mathrm{gt}})\big\|_{2}^{2},$ (21)
 where $\phi_{l}(\cdot)$ denotes the feature maps extracted from the $l$-th VGG-16 layer, and $C_{l},H_{l},W_{l}$ denote the channel count, height, and width of $\phi_{l}$, respectively.
-By comparing representations across selected VGG-16 layers, the perceptual loss captures high-level semantics (e.g., object shapes and textures) and enhances semantic coherence in generated objects.
+By comparing deep feature representations across selected VGG-16 layers, the perceptual loss captures high-level semantic information (e.g., object shapes and textures) and improves semantic coherence in the generated outputs.
 Style Loss.
-The style loss quantifies texture distribution and stylistic characteristics using Gram matrices as in style-transfer methods [77].
+The style loss measures texture and style characteristics via Gram matrices, following style-transfer approaches [79].
 For feature maps $\phi_{l}(x)$ at layer $l$ of VGG-16, the Gram matrix is
@@
 where $\mathcal{L}$ denotes the set of selected layers and $\|\cdot\|_{F}$ the Frobenius norm.
-This statistical matching captures local texture statistics and global stylistic discrepancies, improving the naturalness of restored images and reducing pattern repetition and artifacts.
-Composite Loss.
-The integrated loss combines the above terms.
-The weighted reconstruction term enhances edge clarity and structural continuity in semantic regions; the perceptual term reinforces high-level semantic consistency aligned with human perception; and the style term mitigates texture artifacts across diverse objects and scenes:
+This statistical matching captures both local texture patterns and global style consistency, enhancing realism in restored images while reducing artifacts and repetitive patterns.
+Composite Loss. Finally, we combine these components into a composite loss:
 $\mathcal{L}_{\mathrm{total}}=\lambda_{1}\mathcal{L}_{\mathrm{recon}}+\lambda_{2}\mathcal{L}_{\mathrm{perc}}+\lambda_{3}\mathcal{L}_{\mathrm{style}}\quad(\lambda_{1}=1.0,\,\lambda_{2}=0.1,\,\lambda_{3}=250)$ (24)
-These weights are empirically chosen to balance the differing numerical scales and competing objectives of the loss components during training.
+The reconstruction term sharpens edges in key regions, the perceptual term ensures semantic alignment with human vision, and the style term reduces texture artifacts across varied content.
+These weights balance the varying scales and objectives of the loss components during training.
 4 Experiment
 4.1 Dataset
-Four distinct datasets were employed.
-A custom object-centric teapot dataset (2,000 training, and 200 test samples) constructed in our laboratory featuring $512\times 512$ scene images of multi-view teapot objects against complex backgrounds with artificial items and indoor environments, each paired with corresponding $512\times 512$ ground-truth masks that underwent multi-stage manual calibration to preserve geometric structures and textural details; alongside three COCO2017-derived subsets—“elephant” (2,143 training images, 89 validation as test), “giraffe” (2,546 training, 101 validation as test), and “zebra” (1,916 training, 85 validation as test)—all uniformly resized to $512\times 512$ resolution to ensure comparability.
-A consistent target-aware occlusion protocol was implemented across all datasets, automatically generating masks from target maps to restrict occlusions exclusively to foreground areas, with coverage constrained to $\sim$40% of object regions during both training and inference phases to preserve contextual information while necessitating complex structural reasoning.
+We utilized four distinct datasets for our experiments.
+This included a custom, object-centric teapot dataset (2,000 training and 200 test samples), created in-house.
+It contains $512\times 512$‘ images of teapots from multiple viewpoints, set against complex backgrounds with artificial items and indoor settings.
+Each image has a corresponding $512\times 512$‘ ground-truth mask, meticulously calibrated through a multi-stage manual process to ensure accurate geometric and textural details.
+Additionally, three subsets were derived from COCO2017: “elephant” (2,143 training, 89 validation-test), “giraffe” (2,546 training, 101 validation-test), and “zebra” (1,916 training, 85 validation-test).
+All images were resized to $512\times 512$‘ to maintain consistency.
+A consistent occlusion protocol was applied across all datasets.
+Masks were automatically generated from target maps to ensure occlusions affected only the foreground objects, covering approximately $\sim$40%‘ of their area.
+This balance preserved sufficient context while requiring the model to perform non-trivial structural reasoning during both training and inference.
 4.2 Experimental Setup
 Implementation Details.
-All experiments were conducted on an Ubuntu 20.04 system equipped with an Intel Xeon Platinum 8358P CPU and an NVIDIA GeForce RTX 4090 GPU (24 GB VRAM).
-The framework was implemented in PyTorch 1.10.0 with CUDA 11.3 support.
-Training spanned 50 epochs with a batch size of 4, constrained by available computational resources.
-Optimization employed Adam with a learning rate of $2\times 10^{-4}$, resulting in approximately 0.5 hours of total training time on the specified hardware.
-We adopted a two-stage training protocol: First, the ObjectImagingModule was pretrained exclusively using $\mathcal{L}_{\mathrm{recon}}$ to generate semantically structured intermediate representations.
-This plug-and-play module was subsequently frozen and integrated across all four experimental configurations.
-Second, the Structural Refinement Module (SRM) and Geometric Detail Reconstruction Module (GDRM) were trained using $\mathcal{L}_{\mathrm{perc}}$ and $\mathcal{L}_{\mathrm{style}}$ to enhance semantic coherence and stylistic fidelity.
-This secondary optimization (implemented solely for preliminary validation in Experiment 1) maintained the ObjectImagingModule’s parameters fixed.
-Full implementation code is available at: https://github.com/WYH302/WuYonghao.git.
+All experiments were run on a Ubuntu 20.04 system with an Intel Xeon Platinum 8358P CPU and an NVIDIA GeForce RTX 4090 GPU (24 GB VRAM).
+We implemented our framework in PyTorch 1.10.0 using CUDA 11.3.
+Training lasted for 50 epochs with a batch size of 4, a constraint imposed by our computational resources.
+We used the Adam optimizer with a learning rate of $2\times 10^{-4}$‘.
+The total training time was approximately 0.5 hours on this hardware.
+Our training followed a two-stage protocol.
+First, we pre-trained the ObjectImagingModule using only the reconstruction loss $\mathcal{L}_{\mathrm{recon}}$‘ to produce structured intermediate representations.
+This plug-and-play module was then frozen and integrated into all subsequent experimental setups.
+Second, we trained the Structural Refinement Module (SRM) and Geometric Detail Reconstruction Module (GDRM) with the perceptual loss $\mathcal{L}_{\mathrm{perc}}$‘ and style loss $\mathcal{L}_{\mathrm{style}}$‘ to improve semantic coherence and stylistic fidelity.
+During this second stage, the ObjectImagingModule’s parameters remained fixed.
+This optimization was conducted primarily for preliminary validation in Experiment 1.
+Our full implementation code is available at: https://github.com/WYH302/WuYonghao.git.
 Baseline.
-To rigorously evaluate the efficacy of our proposed method, we benchmark against a representative spectrum of state-of-the-art inpainting approaches selected through a multi-faceted lens.
-Our criteria prioritize comprehensive coverage of methodological evolution, spanning from classical exemplar-based techniques to cutting-edge generative models, while ensuring architectural diversity across diffusion, transformer, and GAN-based paradigms.
-The evaluation suite begins with foundational exemplar-based methods: Criminisi [18], representing classical non-local patch synthesis, establishes a performance baseline for traditional non-deep-learning approaches and highlights inherent limitations in complex structural recovery.
-Its optimized counterpart, PatchMatch [78], provides critical insights into computational efficiency trade-offs through accelerated nearest-neighbor matching.
-Transitioning to modern deep learning, AOT-GAN [79] is included for its attention-driven, long-range dependency modeling via dilated convolutions—directly addressing holistic scene coherence challenges central to our work.
-Edge-Connect [56] offers a strategic comparative lens as a two-stage edge-guided model, enabling explicit evaluation of structural prior utilization against our coarse-to-fine refinement paradigm.
-The hybrid methodology of MISF [80] tests multi-scale feature interaction between traditional and deep components against our biologically inspired framework.
-Finally, contemporary diffusion models LaMa [81] and RePaint [41] are incorporated as cutting-edge references, with LaMa validating spectral processing choices and RePaint contrasting stochastic refinement against our deterministic pipeline.
-All baselines underwent five independent runs with randomized initializations, utilizing official implementations and pre-trained models where available.
+To rigorously evaluate our method, we benchmark against a representative selection of state-of-the-art inpainting approaches.
+Our selection emphasizes methodological diversity, encompassing classical exemplar-based methods to modern generative approaches, including diffusion models, transformers, and GAN architectures.
+We include Criminisi [18] as a foundational exemplar-based method that uses non-local patch synthesis, providing a baseline for traditional approaches while revealing limitations in complex structural recovery.
+PatchMatch [80], an optimized variant, offers insights into computational efficiency through accelerated nearest-neighbor matching.
+From modern deep learning, AOT-GAN [81] is included for its attention-driven modeling of long-range dependencies using dilated convolutions, addressing holistic scene coherence.
+Edge-Connect [44] provides a two-stage edge-guided approach that allows comparison of structural prior usage with our coarse-to-fine framework.
+The hybrid MISF [82] tests multi-scale feature interactions between traditional and deep components, providing a contrast to our bio-inspired approach.
+Finally, we include contemporary diffusion models: LaMa [83] validates our spectral processing choices, while RePaint [41] provides a stochastic refinement contrast to our deterministic approach.
+All baseline methods were executed for five independent runs with random initializations, using official implementations and pre-trained models when available.
 Evaluation metrics.
-We employed four widely used metrics to quantify inpainting performance: (1) SSIM (Structural Similarity Index Measure), which assesses structural similarity between predictions and ground truth (higher is better); (2) PSNR (Peak Signal-to-Noise Ratio), measuring pixel-wise reconstruction fidelity (higher is better); (3) MAE (Mean Absolute Error), which measures average absolute pixel deviation (lower is better); and (4) LPIPS [82], a learned perceptual metric based on deep features (lower is better).
+We used four standard metrics: (1) SSIM measures structural similarity (higher values preferred), (2) PSNR quantifies pixel-level fidelity (higher values better), (3) MAE calculates average pixel deviation (lower values better), and (4) LPIPS [84] evaluates perceptual quality (lower values better).
 4.3 Preliminary theoretical verification
-To isolate and validate the core contribution of our work—the Specific Object-of-Interest Imaging (SIOI) module—we first conducted a controlled experiment using a simplistic subsequent inpainting network (Stage II: SRM+GDRM).
-It is crucial to emphasize that this inpainting network is intentionally simplistic and is not the focus of our innovation.
-Its sole purpose is to serve as a stable and transparent baseline to demonstrate that the high-quality output is primarily attributable to the superior structural and semantic priors generated by the SIOI module (Stage I).
-Unlike conventional image inpainting methods, our model eliminates the need for additional manually annotated corruption masks and is capable of automatically identifying and inpainting the corrupted regions of the target.
-We deliberately retained some black borders to demonstrate that the model does not solely rely on the color black to identify corrupted regions.
+To validate our core contribution—the Specific Object-of-Interest Imaging (SIOI) module—we conducted a controlled experiment using a simplified inpainting network (Stage II: SRM+GDRM).
+This network is intentionally simplified and not our innovation focus.
+Its purpose is to demonstrate that high-quality outputs primarily stem from SIOI’s structural and semantic priors (Stage I).
+Unlike conventional methods, our model automatically identifies and inpaints corrupted regions without manual mask annotation.
+We retained black borders to show the model doesn’t rely solely on color cues for corruption detection.
 As shown in Table 2 and Fig.
-6, even with this simple network, the pipeline achieves exceptional performance on the Teapot dataset.
-This result provides foundational evidence that the SIOI module successfully accomplishes its design goal: to extract a clean, well-defined, and structurally sound representation of the target object from a complex and occluded scene.
-This output serves as an ideal input for anydownstream inpainting task, thereby validating the SIOI module’s efficacy as a standalone component.
+6, this pipeline achieves exceptional performance on the Teapot dataset.
+These results confirm the SIOI module successfully extracts clean, structurally sound object representations from complex, occluded scenes.
+This validates SIOI’s efficacy as a standalone component for downstream inpainting tasks.
 Table 2: Preliminary Theoretical Verification on the teapot dataset (mean $\pm$ std over five runs). ↑ means higher is better, and ↓ means lower is better.
@@
 4.4 Generalized empirical evaluation——Quantitative comparisons
-Having established the standalone effectiveness of the SIOI module, we now evaluate its primary value proposition: its utility as a universal, plug-and-play front-end that can enhance diverse, off-the-shelf inpainting models.
-We integrated our SIOI module with a wide spectrum of state-of-the-art methods, spanning from classical algorithms to modern deep learning approaches.
-To ensure a fair comparison, all baseline methods were re-evaluated using a consistent object-level evaluation protocol, where their outputs were multiplied by corresponding object masks prior to quantitative analysis.
-As summarized in Table 3, the SIOI module consistently enables significant performance improvements across all evaluation metrics.
-The key observation is not the absolute performance of any single model, but the transformative gain achieved by prepending SIOI.
-For instance, SIOI enables Criminisi to achieve a PSNR of 31.552 dB, a 7 dB improvement over the vanilla version.
-More notably, SIOI boosts LaMa to a staggering 45.222 dB PSNR and 0.0023 LPIPS, significantly surpassing all other methods.
-The improvements are especially pronounced on the Teapot dataset, where SIOI+LaMa achieves near-perfect reconstruction (SSIM: 0.9973, LPIPS: 0.0023).These results demonstrate that SIOI is not merely complementary but revolutionary, allowing both traditional and modern algorithms to transcend their inherent limitations by providing a purified, object-centric input.
-Regarding computational efficiency, SIOI acts as a generic adapter, effectively bridging the gap between holistic image processing and efficient object-centric reconstruction.
-Across multiple datasets, models incorporating SIOI consistently achieve superior results in SSIM, PSNR, MAE, and LPIPS metrics compared to those without SIOI.
-Furthermore, models employing SIOI exhibit significantly enhanced perceptual quality.
-These results substantiate the effectiveness of SIOI in achieving high-fidelity reconstruction of objects with complex textures.
-Additionally, the object-centric imaging paradigm facilitates consistent separation of target objects across diverse backgrounds.
-Regarding computational efficiency, SIOI reduces redundant computations by concentrating processing resources on target regions, thereby accelerating inference while maintaining reconstruction quality.
+We next evaluate SIOI’s utility as a universal plug-and-play front-end for enhancing diverse inpainting models.
+We integrated SIOI with state-of-the-art methods spanning classical to modern deep learning approaches.
+For fair comparison, all methods were evaluated using a consistent object-level protocol, with outputs multiplied by corresponding object masks.
+As shown in Table 3, SIOI consistently improves performance across all metrics.
+The key finding is the transformative gain from prepending SIOI, not any single model’s absolute performance.
+For example, SIOI improves Criminisi’s PSNR by 7 dB to 31.552 dB.
+Notably, SIOI boosts LaMa to 45.222 dB PSNR and 0.0023 LPIPS, surpassing all other methods.
+Improvements are most pronounced on the Teapot dataset, with SIOI+LaMa achieving near-perfect reconstruction (SSIM: 0.9973, LPIPS: 0.0023).
+These results show SIOI enables both traditional and modern algorithms to transcend their limitations through purified, object-centric inputs.
+SIOI also improves computational efficiency by bridging holistic image processing and efficient object-centric reconstruction.
+Across datasets, models with SIOI consistently outperform those without it across all metrics.
+Models with SIOI also exhibit enhanced perceptual quality.
+These results confirm SIOI’s effectiveness for high-fidelity reconstruction of complex-textured objects.
+The object-centric paradigm enables consistent target separation across diverse backgrounds.
+Computationally, SIOI reduces redundant processing by focusing on target regions, accelerating inference while maintaining quality.
 Table 3: Experimental Results for Different Datasets and Models, ↑ means higher is better, and ↓ means lower is better.
@@
 RePaint 0.9330 24.037 4.1087 0.1141
-SIOI+Criminisi 0.9841 31.552 0.7396 0.0246
-SIOI+PatchMatch 0.9752 25.912 1.7894 0.0408
-SIOI+Edge-Connect 0.9795 32.378 0.8364 0.0299
-SIOI+AOT-GAN 0.9248 36.940 1.0471 0.0094
-SIOI+MISF 0.9739 33.487 1.0541 0.0674
-SIOI+LaMa 0.9973 45.222 0.1119 0.0023
-SIOI+RePaint 0.9712 28.599 1.6907 0.0645
+SIOI+Criminisi 0.9841(↑4.5%) 31.5520(↑28.3%) 0.7396(↓79.2%) 0.0246(↓74.5%)
+SIOI+PatchMatch 0.9752(↑5.9%) 25.9120(↑22.1%) 1.7894(↓67.2%) 0.0408(↓67.7%)
+SIOI+Edge-Connect 0.9795(↑3.6%) 32.3780(↑29.0%) 0.8364(↓75.2%) 0.0299(↓67.1%)
+SIOI+AOT-GAN 0.9248(↓2.5%) 36.9400(↑45.8%) 1.0471(↓67.8%) 0.0094(↓89.1%)
+SIOI+MISF 0.9739(↑3.9%) 33.4870(↑34.6%) 1.0541(↓71.6%) 0.0674(↓42.8%)
+SIOI+LaMa 0.9973(↑4.7%) 45.2220(↑77.1%) 0.1119(↓96.4%) 0.0023(↓97.2%)
+SIOI+RePaint 0.9712(↑4.1%) 28.5990(↑19.0%) 1.6907(↓58.9%) 0.0645(↓43.5%)
 elephant Criminisi 0.9671 32.615 1.3453 0.0262
@@
 RePaint 0.8624 23.690 5.1769 0.1676
-SIOI+Criminisi 0.9774 33.152 0.8312 0.0190
-SIOI+PatchMatch 0.9721 29.327 1.3736 0.0325
-SIOI+Edge-Connect 0.9636 30.693 1.2479 0.0448
-SIOI+AOT-GAN 0.9341 34.691 1.4350 0.0243
-SIOI+MISF 0.8883 29.070 3.4237 0.1848
-SIOI+LaMa 0.9786 34.706 0.6523 0.0140
-SIOI+RePaint 0.9102 28.956 3.2348 0.1489
+SIOI+Criminisi 0.9774(↑1.1%) 33.1520(↑1.6%) 0.8312(↓38.2%) 0.0190(↓27.5%)
+SIOI+PatchMatch 0.9721(↑1.3%) 29.3270(↑2.4%) 1.3736(↓34.6%) 0.0325(↓26.3%)
+SIOI+Edge-Connect 0.9636(↑0.8%) 30.6930(↓4.0%) 1.2479(↓28.5%) 0.0448(↑5.9%)
+SIOI+AOT-GAN 0.9341(↑11.0%) 34.6910(↑25.5%) 1.4350(↓70.9%) 0.0243(↓60.0%)
+SIOI+MISF 0.8883(↑4.9%) 29.0700(↑22.2%) 3.4237(↓35.3%) 0.1848(↓3.1%)
+SIOI+LaMa 0.9786(↑6.0%) 34.7060(↑33.2%) 0.6523(↓78.0%) 0.0140(↓85.2%)
+SIOI+RePaint 0.9102(↑5.5%) 28.9560(↑22.2%) 3.2348(↓37.5%) 0.1489(↓11.2%)
 giraffe Criminisi 0.9721 29.962 1.5050 0.0184
@@
 RePaint 0.9079 23.563 4.5075 0.1096
-SIOI+Criminisi 0.9788 29.824 1.1000 0.0170
-SIOI+PatchMatch 0.9734 26.327 1.7493 0.0306
-SIOI+Edge-Connect 0.9676 27.527 1.6295 0.0382
-SIOI+AOT-GAN 0.8741 30.861 2.3573 0.0350
-SIOI+MISF 0.9325 27.270 3.0461 0.1215
-SIOI+LaMa 0.9796 31.023 0.9491 0.0129
-SIOI+RePaint 0.9465 27.109 2.8830 0.0992
+SIOI+Criminisi 0.9788(↑0.7%) 29.8240(↓0.5%) 1.1000(↓26.9%) 0.0170(↓7.6%)
+SIOI+PatchMatch 0.9734(↑0.8%) 26.3270(↓4.9%) 1.7493(↓24.0%) 0.0306(↓16.8%)
+SIOI+Edge-Connect 0.9676(↑0.0%) 27.5270(↓8.0%) 1.6295(↓2.7%) 0.0382(↑36.4%)
+SIOI+AOT-GAN 0.8741(↓4.0%) 30.8610(↑15.4%) 2.3573(↓43.2%) 0.0350(↓9.8%)
+SIOI+MISF 0.9325(↑3.6%) 27.2700(↑16.0%) 3.0461(↓34.8%) 0.1215(↑0.6%)
+SIOI+LaMa 0.9796(↑3.9%) 31.0230(↑20.6%) 0.9491(↓63.4%) 0.0129(↓80.7%)
+SIOI+RePaint 0.9465(↑4.3%) 27.1090(↑15.0%) 2.8830(↓36.0%) 0.0992(↓9.5%)
 zebra Criminisi 0.9677 27.598 2.2333 0.0236
@@
 RePaint 0.8677 21.544 8.0205 0.1499
-SIOI+Criminisi 0.9750 27.389 1.6940 0.0211
-SIOI+PatchMatch 0.9675 24.691 2.4553 0.0365
-SIOI+Edge-Connect 0.9632 26.089 2.2360 0.0387
-SIOI+AOT-GAN 0.9070 28.025 3.2725 0.0527
-SIOI+MISF 0.8960 24.048 6.2437 0.1556
-SIOI+LaMa 0.9772 28.811 1.3896 0.0144
-SIOI+RePaint 0.9130 24.309 5.9439 0.1365
+SIOI+Criminisi 0.9750(↑0.8%) 27.3890(↓0.8%) 1.6940(↓24.1%) 0.0211(↓10.6%)
+SIOI+PatchMatch 0.9675(↑4.5%) 24.6910(↑9.1%) 2.4553(↓40.8%) 0.0365(↓57.4%)
+SIOI+Edge-Connect 0.9632(↑0.5%) 26.0890(↓6.5%) 2.2360(↓28.0%) 0.0387(↑47.1%)
+SIOI+AOT-GAN 0.9070(↑7.3%) 28.0250(↑25.1%) 3.2725(↓66.0%) 0.0527(↑8.4%)
+SIOI+MISF 0.8960(↑4.4%) 24.0480(↑12.3%) 6.2437(↓24.4%) 0.1556(↓2.5%)
+SIOI+LaMa 0.9772(↑4.7%) 28.8110(↑18.6%) 1.3896(↓62.9%) 0.0144(↓78.9%)
+SIOI+RePaint 0.9130(↑5.2%) 24.3090(↑12.8%) 5.9439(↓25.9%) 0.1365(↓8.9%)
 4.5 Generalized empirical evaluation——Qualitative comparisons
-Qualitative results further corroborate the role of SIOI as a powerful pre-processing stage.
-By providing a clean, background-suppressed, and structurally coherent representation of the target object, SIOI enables even conventional algorithms to produce more semantically consistent and visually plausible results.
-Fig.
-7, Fig.
-8, Fig.
-9, Fig.
-10, Fig.
-11, Fig.
-12, Fig.
-13, Fig.
-14, presents qualitative results that highlight characteristic behaviors and failure modes of the benchmark models.
-Criminisi [18] fundamentally lacks semantic understanding in its non-local patch matching, leading to severe structural discontinuities and semantically inconsistent textures in complex scenes.
-Similarly, PatchMatch [78] remains constrained by hand-crafted features and low-level metrics, failing to capture high-level semantics or global coherence, which results in visible patch artifacts—a byproduct of its localized/static patch-level processing that becomes evident under gradient-rich lighting and on non-rigid surfaces.
-AOT-GAN [79] exhibits jagged artifacts and discontinuous textures in complex regions, these failures stem from its use of dilated convolutions for feature aggregation.
-Edge-Connect [56], while effective in many cases, suffers from visible spatial misalignment due to its over-reliance on imperfect Canny edges, which do not correspond to perceptually relevant structures and disrupt spatial coherence in repetitive textures (e.g., grass).
-MISF [80] produces overly smooth, artificial-looking edges and illumination inconsistencies, resulting from its multi-scale regularization and strong smoothing priors as well as incompatible traditional and deep feature interactions.
-LaMa [81] over-smooths detailed regions and produces blurry or insufficiently sharp outputs despite its efficient Fourier-based design.
-RePaint [41] often yields overly smooth or blurry outputs with incongruous textures due to its slow.
-Without using our method, existing methods exhibit characteristic failures due to background interference and a lack of object-level understanding, resulting in color and texture details related to the background that are completely inconsistent with their original objects.
-In contrast, the integration of the SIOI module consistently guides the inpainting focus, resulting in improved edge sharpness, texture consistency, and semantic coherence, regardless of the backend inpainting algorithm used.
-Experimental results across multiple object datasets show that our approach performs particularly well on rigid objects (e.g., teapots), while achieving relatively weaker results on non-rigid objects such as giraffes, zebras, and elephants.
-We attribute this difference to the higher morphological variability and structural complexity of non-rigid categories, which may require larger or more adaptive model parameters to adequately capture their semantic and textural characteristics.
-Preliminary observations suggest a potential linear relationship among semantic complexity, training data scale, and model parameter quantity—a phenomenon that merits further quantitative investigation in future work.
+Qualitative results confirm SIOI’s effectiveness as a preprocessing stage.
+SIOI provides clean, background-suppressed object representations, enabling conventional algorithms to produce more consistent and plausible results.
+As shown in Figures 7-14, benchmark models exhibit characteristic failure modes: Criminisi [18] lacks semantic understanding, causing structural discontinuities and inconsistent textures; PatchMatch [80] relies on hand-crafted features, producing visible artifacts due to localized patch processing; AOT-GAN [81] shows jagged artifacts from dilated convolutions; Edge-Connect [44] suffers spatial misalignment from over-reliance on Canny edges; MISF [82] produces oversmoothed edges and illumination inconsistencies; LaMa [83] oversmooths details, producing blurry outputs; RePaint [41] yields blurry outputs with incongruous textures.
+Without SIOI, methods fail due to background interference and lack of object understanding.
+With SIOI, all methods show improved edge sharpness, texture consistency, and semantic coherence.
+Our approach excels on rigid objects (e.g., teapots) but shows relatively weaker results on non-rigid objects (giraffes, zebras, elephants).
+This difference stems from higher morphological variability in non-rigid categories, requiring larger or more adaptive models.
+This suggests a relationship between semantic complexity, data scale, and model size, warranting future study.
 Figure 7: Qualitative comparison of exemplar inpainting results (teapot dataset).
@@
 In fact, a detailed ablation experiment has already been conducted above, So this is just a simple supplement to the Preliminary theoretical verification experiment, also conducted on the Teapot dataset.
-This experiment is based on the preliminary theoretical verification experiment.
-To evaluate the effectiveness of our methodology, we conducted ablation studies under identical experimental conditions.
-Key components were systematically removed, and the resulting imaging outcomes were comparatively analyzed to elucidate performance differences.
-As shown in Fig.
+We conducted additional ablation studies on the Teapot dataset under identical conditions.
+Key components were systematically removed to analyze performance differences, as shown in Fig.
 15.
 Figure 15: Illustration of ablation experiment.
-Ablation I: In Ablation I, the state I network was removed while retaining the object-centric input–output pipeline.
-The corresponding results are presented in Tab.
-4.
-The resulting model showed marginal superiority in certain quantitative metrics compared to the proposed method—SSIM (0.976 vs.
-0.978), PSNR (33.95 dB vs.
-33.86 dB), MAE (1.313 vs.
-1.605), and LPIPS (0.017 vs.
-0.018)—but these differences were statistically insignificant in terms of perceptual quality.
-However, subjective evaluation revealed compromised edge localization, manifested as jagged artifacts and global blurring.
-This degradation confirms the critical role of Stage I in achieving high structural precision.
-Furthermore, visual comparisons clearly demonstrated the superior inpainting quality achieved when Stage I was integrated.
-The marginal improvements observed in PSNR and MAE, despite perceptual and qualitative degradation, further highlight that such metrics can be misleading indicators of inpainting quality, as they from the loss function’s direct penalization of RGB deviations, do not consistently align with human perceptual judgment.
-Ablation II: In Ablation II, the object-centric I/O was replaced with full-scene inputs.
-The model processed the entire image, including the extensive and complex background, which resulted in severe performance degradation.
-This represents the standard paradigm used by most existing methods, which our SIOI approach aims to supersede for object-centric tasks.
-Quantitative metrics showed substantial deterioration: SSIM: 0.964, PSNR: 29.43 dB, MAE: 2.085, LPIPS: 0.065.
-Visual inspection confirmed that this configuration yielded the poorest inpainting quality among all compared models.
-It exhibited pronounced detail loss, extensive artifacts throughout the restored regions, and markedly impaired image clarity.
-Furthermore, the model’s inability to effectively extract features, particularly from the edges of degraded regions, resulted in insufficient feature representation during inpainting.
-This shortfall directly led to blurry reconstruction and pervasive artifact generation.
-This ablation highlights the fundamental advantage of our object-first paradigm: it dramatically reduces the solution space by focusing computation exclusively on the target, thereby avoiding the pitfalls of whole-image processing.
-In contrast to both ablated versions, the full model (with SIOI) achieved superior performance, delivering sharp edges, semantically coherent structures, and photorealistic textures.
-This provides compelling evidence that our two-stage pipeline, headed by the SIOI module, effectively addresses the bottlenecks of traditional inpainting by first resolving the ”where and what” of the target object before addressing the ”how” of its inpainting.
-The SIOI module acts as a powerful pre-processor that can be plugged into any inpainting network to elevate its performance on object-centric tasks.
-Table 4: Comparison of ablation study results on the Teapot dataset. (mean $\pm$ std over five runs). ↑ means higher is better, and ↓ means lower is better.
+Ablation I: Removing Stage I while retaining object-centric I/O (Tab.
+4) showed marginal metric improvements (SSIM: 0.976 vs.
+0.978, PSNR: 33.95 dB vs.
+33.86 dB, MAE: 1.313 vs.
+1.605, LPIPS: 0.017 vs.
+0.018), but these were statistically insignificant.
+Subjectively, edge localization degraded with jagged artifacts and blurring, confirming Stage I’s critical role for structural precision.
+The metric improvements, despite perceptual degradation, highlight that pixel-based metrics can mislead due to their focus on RGB deviations rather than human perception.
+Ablation II: Replacing object-centric I/O with full-scene inputs (standard paradigm) caused severe performance degradation: SSIM: 0.964, PSNR: 29.43 dB, MAE: 2.085, LPIPS: 0.065.
+Visual inspection showed poorest quality with detail loss, artifacts, and impaired clarity.
+The model’s inability to extract features from degraded regions led to blurry reconstruction and artifacts.
+This highlights our object-first advantage: focusing computation on the target avoids whole-image pitfalls.
+In contrast, the full model (with SIOI) achieved superior performance with sharp edges, semantic coherence, and photorealistic textures.
+This confirms that our two-stage pipeline effectively addresses traditional inpainting bottlenecks by first resolving target object localization and identity before inpainting.
+SIOI serves as a powerful pre-processor that enhances any inpainting network for object-centric tasks.
+Table 4: Comparison of ablation study results on the Teapot dataset. (mean $\pm$ std over five runs). ↑ means higher is better, and ↓ means lower is better (teapot dataset).
 […]
@@
 4.7 Robustness Analysis
-This experiment is based on the preliminary theoretical verification experiment, also conducted on the Teapot dataset.
-To evaluate the generalization ability of our method under challenging conditions, we conducted robustness experiments on the following scenarios:
+This study extends the preliminary theoretical verification experiment on the Teapot dataset. To assess generalization under challenging conditions, we tested robustness in scenarios:
   • 1.
@@
     Motion blur: Linear kernel with $15\times 15$ pixels
-It should be noted that these challenging scenarios were not part of the training data.
-For qualitative assessment, two representative samples were selected per category.
-Quantitative evaluation was omitted due to the limited number of available samples for each specific condition, as illustrated in Fig.
+Note that these scenarios were absent from training data.
+For qualitative assessment, we selected two samples per category.
+Quantitative evaluation was omitted due to limited samples per condition, as shown in Fig.
 16.
-The qualitative results demonstrate the SIOI module’s notable robustness in several challenging conditions.
-It consistently maintained satisfactory target extraction performance under low illumination, small targets, salt-and-pepper noise, and motion blur.
-This resilience can be attributed to the module’s design principles: its spectral processing (FFC) is inherently robust to global illumination changes and motion artifacts, while its attentional mechanisms, inspired by V4 cortical processing, enable it to amplify and track small or salient targets despite noise.
-Performance degradation emerged in scenarios with three or more coexisting objects, as the competition for attentional resources increased.
-This observation directly mirrors the biological phenomenon of attentional limitation described in Desimone and Duncan’s biased competition model [66], validating the bio-inspired nature of our approach even in its failure modes.
-The most substantial degradation occurred under high-power Gaussian noise.
-This is a recognized challenge for feature-based attentional systems both in biological and artificial networks.
-The SIOI module’s reliance on gradient and frequency information for target selection makes it particularly susceptible to this type of non-structured, high-entropy noise, which overwhelms the salient features the module is designed to detect.
-This result precisely highlights the critical role and inherent vulnerability of the front-end SIOI module: its output quality is the foundational prior for all subsequent processing.
-This underscores the importance of our modular design—the failure is contained and diagnosed at the SIOI stage, preventing a cascading collapse of the entire system and clearly pointing to the component that requires improvement.
-These findings provide a clear roadmap for evolving the SIOI module.
-Future work will focus on enhancing its robustness against unstructured noise, potentially by incorporating neuromorphic noise-adaptation mechanisms inspired by the auditory brainstem or retinal adaptive gain control.
-This will further solidify the SIOI module’s role as a robust, plug-and-play pre-processor for object-centric vision tasks.
-In summary, the robustness analysis confirms that the SIOI module provides a reliable and highly effective object-specific prior across a wide spectrum of challenging conditions.
-Its performance profile—strong in structured degradation but susceptible to unstructured noise—precisely characterizes its operational envelope and offers valuable insights for its application and further development as a general-purpose biological vision-inspired front end.
-Figure 16: Illustration of robustness experiment.
+Qualitative results show SIOI’s robustness in several challenging conditions.
+It maintained satisfactory performance under low illumination, small targets, salt-and-pepper noise, and motion blur.
+This resilience stems from its design: spectral processing (FFC) resists global illumination changes and motion artifacts, while attentional mechanisms, inspired by V4, amplify and track small targets despite noise.
+Performance degraded with three or more objects due to attentional resource competition.
+This observation directly mirrors the biological phenomenon of attentional limitation described in Desimone and Duncan’s biased competition model [68].
+Severe degradation occurred under high-power Gaussian noise.
+This challenge is known for feature-based attentional systems in biological and artificial networks.
+SIOI’s reliance on gradient and frequency information makes it susceptible to non-structured, high-entropy noise, overwhelming salient features.
+This highlights SIOI’s critical role and shortcoming: its output quality underpins all subsequent processing.
+Our modular design should contains failures at SIOI stage, preventing system collapse and pinpointing improvements.
+These findings guide SIOI’s evolution.
+Future work will enhance robustness against unstructured noise using neuromorphic adaptations from auditory brainstem or retinal gain control.
+This will solidify SIOI’s role as a plug-and-play pre-processor for object-centric vision.
+In summary, SIOI provides reliable object-specific priors across challenging conditions.
+Its profile—strong against structured degradation but susceptible to unstructured noise—defines its operational envelope and guides application as a bio-inspired front end.
+Figure 16: Illustration of robustness experiment (teapot dataset).
 5 Discussion
-Image impainting is a job without a standard answer and often leads to controversy, as shown in Fig.
+Image inpainting lacks a definitive standard solution and frequently yields contentious results, as illustrated in Fig.
 17.
-We don’t even know if the black area should be an buffalo head or a zebra head.
-The same model, when the same image is input multiple times, will yield different results.
-And the work proposed in this article can solve this problem.
-This work introduces Specific Object-of-Interest Imaging (SIOI)—a novel, bio-inspired paradigm that fundamentally rethinks the pipeline for object-centric visual computing tasks.
-Departing from conventional holistic processing, SIOI provides a plug-and-play front-end module that first isolates and enhances the target object, generating a refined structural and semantic prior that can be seamlessly integrated into any subsequent processing network.
-Our experimental validation demonstrates that incorporating the SIOI module consistently elevates the performance of diverse inpainting architectures.
-This is not a contest between inpainting models; rather, it is a demonstration that providing a high-quality, object-centric prior is a more impactful lever for improvement than designing yet another holistic inpainting network.
-Figure 17: Misunderstand of image impainting.
-This approach aligns with the ventral pathway’s ”object-before-background” processing strategy, where attentional mechanisms in areas like V4 selectively amplify target features while suppressing irrelevant background clutter[83].
-Our SIOI module computationally instantiates this principle, thereby mitigating the fundamental issue of information redundancy that plagues whole-image methods.
-Compared to GAN- or diffusion-based approaches that operate on the entire scene, our method avoids propagating errors from background regions by focusing computation on the target, akin to V4’s role in contour integration and feature binding.
-The robustness of the SIOI module under extreme conditions (e.g., low illumination, noise) further underscores its utility as a reliable pre-processor.
+It is often ambiguous whether a missing region should be inpainted as one object or another (e.g., a buffalo or zebra head).
+Furthermore, the same model can produce varying outputs for an identical input across different runs.
+Our proposed framework addresses this challenge by introducing the Specific Object-of-Interest Imaging (SIOI) paradigm—a bio-inspired approach that reconceptualizes the processing pipeline for object-centric vision tasks.
+Departing from conventional holistic processing, SIOI provides a plug-and-play front-end module that first isolates and enhances the target object, generating a refined structural and semantic prior for any subsequent processing network.
+Our experiments demonstrate that incorporating SIOI consistently elevates the performance of diverse inpainting architectures.
+Thus, our contribution is not to propose a superior inpainting network, but to demonstrate that providing a high-quality, object-centric prior is a more impactful strategy than developing another holistic model.
+Figure 17: Illustration of challenges in image inpainting.
+This approach mirrors the ventral visual pathway’s ”object-before-background” processing strategy, where attentional mechanisms (e.g., in area V4) enhance relevant features while suppressing background clutter [85].
+Our SIOI module computationally implements this principle, mitigating the core issue of informational redundancy inherent in whole-image processing.
+Unlike GAN- or diffusion-based methods that process the entire scene, our approach avoids error propagation from background regions by concentrating computation on the target object, analogous to the role of V4 in contour integration and feature binding.
+The robustness of SIOI under challenging conditions (e.g., low illumination, noise) further highlights its utility as a reliable pre-processing module.
 Its behavior is consistent with the object-based attention mechanism in biological vision, which dynamically modulates early visual areas (V1-V4) to maintain target salience amidst distraction.
-It is important to note that the performance of any subsequent model is contingent on the quality of the prior provided by SIOI.
-This dependency mirrors the dorsal-ventral synergy in the brain, where spatial localization guides object recognition.
+Notably, the performance of any downstream model is contingent on the quality of the prior provided by SIOI.
+This dependency parallels the dorsal-ventral synergy in the brain, wherein spatial localization guides object recognition.
 This highlights that advancing object-level pre-processing is a critical frontier in itself.
-We emphasize that the inpainting network presented in Stage II serves primarily as a proof-of-concept vehicle to validate the efficacy of the SIOI prior.
-our architecture may not represent the ultimate performance ceiling, and that future researchers may develop superior networks.
+We emphasize that the inpainting network in Stage II serves primarily as a proof-of-concept to validate the efficacy of the SIOI prior.
+Consequently, our architecture does not represent the performance ceiling, and future work may develop superior networks.
 The paramount innovation of this work is not this specific network, but the SIOI paradigm itself.
 This paradigm shift—from direct holistic inpainting to object-first prior generation—is our central contribution.
-The generalization of the SIOI module is inherently tied to the semantic definition of a ”specific object.” This is not merely a limitation but a defining characteristic and a research opportunity.
-But we must frankly admit that the generalization of this technology is not strong due to its focus on specific objects.
-And each semantic object represents a type of dataset, which also requires retraining the object imaging network.
-For example, whether cat is an object, or is it divided into jungle cat felis chaus, sand cat felis Margarita, black foot cat felis nigripes, and wild cat felis silvestris(which is also more refined as a specific object) needs further research.
-The module’s performance can be scaled and specialized, offering a practical pathway for technology transfer, allowing the SIOI paradigm to be precisely tailored to application-specific needs.
-Future work will explore integrating prefrontal-cortex-like task-driven modulation to enable dynamic reconfiguration of the SIOI module for different object categories and end goals, advancing toward truly adaptive and biologically plausible computational vision systems.
+The generalization capability of the SIOI module is inherently linked to the semantic definition of a ”specific object.” This is not merely a limitation but a defining characteristic and a research opportunity.
+We acknowledge that the current generalization capability is limited due to its focus on specific object categories.
+Each semantically distinct object category effectively constitutes a separate dataset, necessitating retraining of the object imaging network.
+The granularity of this definition—for instance, whether ”cat” constitutes a single category or should be divided into subspecies (e.g., Felis chaus, Felis margarita)—requires further investigation.
+However, the module’s performance can be scaled and specialized, providing a practical pathway for technology transfer and enabling precise tailoring to application-specific needs.
+Future work will explore integrating prefrontal-cortex-like, task-driven modulation to enable dynamic reconfiguration of SIOI for different object categories and goals, advancing toward adaptive and biologically plausible vision systems.
 6 Conclusion
-In conclusion, this work proposes Specific Object-of-Interest Imaging (SIOI) as a novel, bio-inspired pre-processing paradigm for object-centric visual tasks.
-SIOI operates by first isolating and enhancing the target object, providing a high-fidelity prior that dramatically improves the performance of subsequent vision modules.
-Extensive experiments confirm that the SIOI module is a powerful, plug-and-play front-end that consistently improves various inpainting models by suppressing background interference and providing a coherent object-centric prior.
-Its design is grounded in the ”object-first” processing hierarchy of the ventral visual pathway, offering a computational solution to the problem of information redundancy.The framework demonstrates strong robustness across challenging conditions, reflecting the resilience of biological attentional systems.
-Limitations arise from reliance on initial object generation, which may impair boundary guidance.
-We envision the SIOI paradigm extending beyond inpainting to benefit a wide array of object-centric applications, including segmentation, detection, and video tracking, wherever suppressing background clutter and focusing computational resources on the target is paramount.
-Future work will explore task-driven modulation and fMRI validation, advancing toward biologically adaptive imaging and broader applications in neuroimaging and perceptual assistance.
+In conclusion, we propose Specific Object-of-Interest Imaging (SIOI), a bio-inspired preprocessing paradigm for object-centric visual tasks.
+It operates by isolating and enhancing the target object, providing a high-fidelity prior that enhances subsequent vision modules.
+Extensive experiments confirm SIOI as a powerful, plug-and-play front-end that improves diverse inpainting models by suppressing background interference and providing a coherent object-centric prior.
+Grounded in the ventral pathway’s ”object-first” hierarchy, it offers a computational solution to information redundancy.
+The framework demonstrates robustness under challenging conditions, reflecting the resilience of biological attention.
+A limitation is its reliance on initial object generation, which can impair boundary guidance.
+The SIOI paradigm can benefit other object-centric tasks like segmentation, detection, and tracking, where suppressing background clutter is critical.
+Future work includes task-driven modulation and fMRI validation, advancing toward biologically adaptive imaging for neuroimaging and perceptual assistance.
 Table 5: Extended Mathematical Notation Guide with Biological Correlations